### v1

# TELEClass: Taxonomy Enrichment and LLM-Enhanced Hierarchical Text Classification with Minimal Supervision

Abstract.
Hierarchical text classification aims to categorize each document into a set of classes in a label taxonomy. Most earlier works focus on fully or semi-supervised methods that require a large amount of human annotated data which is costly and time-consuming to acquire. To alleviate human efforts, in this paper, we work on hierarchical text classification with the minimal amount of supervision: using the sole class name of each node as the only supervision. Recently, large language models (LLM) show competitive performance on various tasks through zero-shot prompting, but this method performs poorly in the hierarchical setting, because it is ineffective to include the large and structured label space in a prompt. On the other hand, previous weakly-supervised hierarchical text classification methods only utilize the raw taxonomy skeleton and ignore the rich information hidden in the text corpus that can serve as additional class-indicative features. To tackle the above challenges, we propose TELEClass, Taxonomy Enrichment and LLM-Enhanced weakly-supervised hierarchical text Classification, which (1) automatically enriches the label taxonomy with class-indicative topical terms mined from the corpus to facilitate classifier training and (2) utilizes LLMs for both data annotation and creation tailored for the hierarchical label space. Experiments show that TELEClass can outperform previous weakly-supervised hierarchical text classification methods and LLM-based zero-shot prompting methods on two public datasets.
Weakly-Supervised Text Classification, Hierarchical Text Classification, Taxonomy Enrichment, Large Language Model
$*{}^{*}$qual Contribution.
1. Introduction
Hierarchical text classification, aiming to classify documents into one or multiple classes in a label taxonomy, is an essential and fundamental task in text mining and natural language processing. Compared with standard text classification where label space is flat and relatively small (e.g., less than 20 classes), hierarchical text classification is more challenging given the larger and more structured label space and the existence of fine-grained and long-tail classes in the taxonomy. Most earlier works tackle this task in fully supervised (You et al., 2019; Huang et al., 2019; Chen et al., 2021) or semi-supervised settings (Berthelot et al., 2019; Gururangan et al., 2019), and different models are proposed to learn from a substantial amount of human-labeled data. However, acquiring human annotation is often costly, time-consuming, and not scalable.
Figure 1. An example document tagged with 3 classes. We automatically enrich each node with class-indicative terms and utilize LLMs to facilitate classification.
Recently, large language models (LLM) such as GPT-4 (OpenAI, 2023) and Llama2 (Touvron et al., 2023) have demonstrated strong performance in flat text classification through zero-shot or few-shot prompting (Sun et al., 2023). However, applying LLMs in hierarchical settings, with large and structured label spaces, remains challenging. Directly including hundreds of classes in prompts is ineffective, leading to structural information loss, increased computational costs, and diminished clarity for LLMs in focusing on critical information. Along another line of research, Meng et al. (2019) propose to train a moderate-size text classifier by utilizing a small set of keywords or labeled documents for each class. However, compiling keyword lists for hundreds of classes and obtaining representative documents for each specific and niche category still demand significant human efforts. Shen et al. (2021) study the hierarchical text classification with minimal supervision, which takes the class name as the only supervision signal. Specifically, they introduce TaxoClass, a method that generates pseudo-labels with a textual entailment model for classifier training. However, this method overlooks additional class-relevant features in the text and suffers from unreliable pseudo-label selection due to the entailment model’s domain shift.
In this study, we advance minimally supervised hierarchical text classification by integrating corpus-based taxonomy enrichment and leveraging LLMs tailored for the hierarchical label structure. First, we enrich the existing label taxonomy skeleton with automatically extracted class-specific topical terms from the corpus, enhancing supervision signals and thus improving the pseudo-label quality for classifier training. Second, we explore the application of LLMs in hierarchical text classification from two perspectives. Specifically, we enhance LLM annotation efficiency and effectiveness through a taxonomy-guided candidate search and optimize LLM-generated document accuracy by using taxonomy paths to create more precise pseudo-data.
Leveraging the above ideas, we introduce TELEClass: Taxonomy Enrichment and LLM-Enhanced weakly-supervised hierarchical text Classification. TELEClass consists of four major steps: (1) LLM-Enhanced Core Class Annotation, where we identify document “core classes” (i.e., fine-grained classes that most accurately describe the documents) by first using a textual entailment model for top-down candidate search on the taxonomy and then prompting an LLM to select the most precise core classes. (2) Corpus-Based Taxonomy Enrichment, where we analyze the taxonomy structure and the roughly classified documents to identify class-indicative topical terms through semantic and statistical analysis, enriching the label taxonomy with more features. For example, the “conditioner” class in Figure 1 is enriched with key terms like “moisture” and “soft hair”, which distinguish it from its sibling classes. (3) Core Class Refinement with Enriched Taxonomy, where we embed documents and classes based on the enriched label taxonomy and refine the initially selected core classes by identifying the most similar classes for each document. (4) Text Classifier Training with Path-Based Data Augmentation, where we sample label paths from the taxonomy and guide the LLM to generate pseudo documents most accurately describing these fine-grained classes. Finally, we train the text classifier on two types of pseudo-labels, the core classes and the generated data, with a simple text-matching network and multi-label training strategy.
The contributions of this paper are summarized as follows: (1) We propose TELEClass, a new method for minimally supervised hierarchical text classification, which requires only the class names of the label taxonomy as supervision to train a multi-label text classifier. (2) We propose to enrich the label taxonomy with class-indicative topical terms mined from the text corpus, based on which we utilize an embedding-based document-class matching method to improve the pseudo-label quality. (3) We study two ways of adopting large language models to the hierarchical text classification setting, which can improve the pseudo-label quality and solve the data scarcity issue for fine-grained classes. (4) Extensive experiments on two datasets show that TELEClass can outperform strong weakly-supervised hierarchical text classification methods and zero-shot LLM prompting methods.  [^1] Code and datasets are available upon request now and will later be available online.
2. Problem Definition
Following Shen et al. (2021), we define the minimally-supervised hierarchical text classification task as to train a text classifier that can categorize each document into multiple nodes on a label taxonomy by using the name of each node as the only supervision. For example, in Figure 1, the input document is classified as “hair care”, “shampoo”, and “scalp treatment”.
[…]
3. Methodology
In this section, we will introduce TELEClass consisting of the following modules: (1) LLM-enhanced core class annotation, (2) corpus-based taxonomy enrichment, (3) core class refinement with enriched taxonomy, and (4) text classifier training with path-based data augmentation. Figure 2 shows an overview of TELEClass.
Figure 2. Overview of the TELEClass framework.
3.1. LLM-enhanced Core Class Annotation
Inspired by previous studies, we first tag each document with its “core classes”, which are defined as a set of classes that can describe the document most accurately (Shen et al., 2021). This process also mimics the process of human performing hierarchical text classification: first, select a set of most essential classes for the document and then trace back to their ancestors to complete the labeling. For example, in Figure 1, by first tagging the document with “shampoo” and “scalp treatment”, we can easily find its complete set of classes.
To mine the core classes for each document, existing methods utilize a textual entailment model (e.g., RoBERTa-large-MNLI) and combine tree-based top-down candidate searching with corpus-level statistical analysis for core class selection. Specifically, the similarity score between a document and a class is defined as the entailment score from the document to a class-specific hypothesis such as “This example is [Class Name]”. These similarity scores facilitate a top-down search in the taxonomy to shortlist candidate classes, which are then evaluated across the corpus to determine the most reliable core classes.
However, there are two limitations of this method. First, the textual entailment model, not being directly trained for document classification, offers inconsistent zero-shot performance, particularly in distinguishing closely related, fine-grained classes within a hierarchical structure. Second, given a pair of premise (document) and hypothesis, the entailment model is trained to predict a distribution over three classes, namely entailment, neutral, and contradiction. Therefore, it is questionable whether the generated scores can support corpus-level comparison, because the scores between different pairs may not be directly comparable. Both limitations will lead to noisy core class selection and adversely impact the classifier performance.
In this work, we propose to utilize an LLM to enhance the core class annotation step. These LLMs are trained on very large textual data and human preference data and shown to have superior zero-shot ability for the document classification task with a small and flat label space (Sun et al., 2023). However, LLMs cannot effectively understand the hierarchical label space, which is large and structured and cannot be easily represented in a prompt. Therefore, we precede the LLM annotation with the top-down candidate search method using a textual entailment model.
Specifically, for each document $d_{i}$, we start from the Root node at level 0 of the taxonomy by adding it into a queue. For each class at level $l$ in the queue, we select $l+3$ of its children that have the highest similarity to document $d_{i}$, where the similarity score $sim^{\text{entail}}(c,d_{i})$ is obtained from a textual entailment model. Then, among all the selected children at level $l+1$, we keep the top $(l+2)^{2}$ of them with the highest path-based similarity scores and add to the queue, where the path-based similarity $sim^{\text{path}}(c,d_{i})$ is defined recursively as,
$sim^{\text{path}}(\text{Root},d)$ $=1,$
$sim^{\text{path}}(c,d)$ $=\max_{c_{p}\in Par(c)}sim^{\text{path}}(c_{p},d_{i})\cdot sim^{\text{entail}}(c,d_{i}),$
where $Par(c)$ is the set of all parents of $c$. This process finishes when there is no class in the queue, and all the nodes (other than Root) ever pushed into the queue are treated as candidate core classes for document $d_{i}$. To use an LLM to annotate core classes, we first construct an instruction for each document, which asks the LLM to select the most accurate classes for the document from its candidate core classes. Then, the document is fed to the LLM as a query and it will generate a list of core classes according to the document and candidate core classes.
Our LLM-enhanced core class annotation step will assign an initial set of core classes (denoted as $\mathbb{C}^{0}_{i}$) for each document $d_{i}\in\mathcal{D}$. In the next two steps, we introduce how TELEClass leverages the initial core classes to enrich the label taxonomy with class-indicative features. This enriched taxonomy is subsequently employed to refine the initial core classes, utilizing embedding-based similarity scores that are more comparable across documents.
3.2. Corpus-Based Taxonomy Enrichment
Due to the hierarchical nature of taxonomy, fined-grained classes at lower levels are often hard to distinguish with only class label names. Therefore, we propose to enrich the weak supervision, i.e. the label taxonomy, with a set of class-indicative topical terms for each class. These enriched terms can help better interpret the class and thus facilitate text classification. For example, in Figure 2, “shampoo” and “conditioner” are two fine-grained classes that are similar to each other. We can effectively separate the two classes by identifying a set of class-specific terms such as “flakes” for “shampoo” and “moisture” for “conditioner”.
In this work, we combine statistical document occurrence features and semantic similarity and further contrast a class with its siblings to identify a set of key terms that can distinguish it from its most similar nodes in the taxonomy. Notice that, because the input taxonomy $\mathcal{T}$ is a directed acyclic graph rather than a simple tree, a node $c$ can have multiple sets of siblings, each of which corresponds to one of $c$’s parent nodes. Besides, because we have already tagged each document with its core classes, we can use them as rough classification results and form a document cluster for each class to help the enrichment step.
Formally, given a class $c\in\mathcal{C}$ and its siblings corresponding to one of its parents $c_{p}$, $Sib(c,c_{p})=\{c^{\prime}\in\mathcal{C}|\langle c_{p},c^{\prime}\rangle\in\mathcal{R}\}$, $\mathbf{c}_{p}\in Par(c)$, we find a set of class-indicative topical terms of $c$ corresponding to $c_{p}$, denoted as $T(c,c_{p})=\{t_{1},t_{2},\dots,t_{k}\}$. Each term in $T(c,c_{p})$ can signify the class $c$ and distinguish it from its siblings under $c_{p}$. To identify these class-indicative terms, we construct a set of documents $D^{0}_{c}$ for each class $c$, containing all the documents whose initial core classes contain $c$ or its descendants,
(1) $\small D^{0}_{c}=\left\{d_{i}\in D|\exists c^{\prime}\in\mathbb{C}^{0}_{i},c^{\prime}\in\{c\}\cup Des(c)\right\},$
where $Des(c)$ is the set of all descedants of $c$. We consider the following 3 factors for class-indicative term selection: popularity, distinctiveness, and semantic similarity.
Popularity. First, a class-indicative term of $c$ should be frequently mentioned by the documents about $c$, denoted as its popularity within $c$. We calculate the popularity of a term $t$ within a class $c$ as,
(2) $\small pop(t,c)=\log(1+df(t,D^{0}_{c})),$
where the document frequency $df(t,D)$ stands for the number of documents in $D$ that mention $t$.
Distinctiveness. Second, a class-indicative term $t$ for a class $c$ should also be infrequent in its siblings so that a document containing $t$ likely indicates that it is about $c$ instead of $c$’s siblings. We define the distinctiveness score of a term $t$ for a class $c$, corresponding to its siblings under $c_{p}$, as,
(3) $\small dist(t,c,c_{p})=\frac{\exp(BM25(t,D^{0}_{c}))}{1+\sum_{c^{\prime}\in Sib(c,c_{p})}\exp(BM25(t,D^{0}_{c^{\prime}}))},$
where $BM25(\cdot,\cdot)$ denotes the BM25 relevance function (Robertson and Zaragoza, 2009).
[…]
Finally, we define the affinity score between a term $t$ and a class $c$ corresponding to parent $p$ to be the geometric mean of the above scores, with a weight parameter $\alpha$,
(4) $\small aff(t,c,c_{p})=pop(t,c)^{\alpha}\cdot dist(t,c,c_{p})^{1-\alpha}\cdot\cos(\mathbf{t},\mathbf{c}).$
Taxonomy Enrichment. We first apply a phrase mining tool, AutoPhrase (Shang et al., 2018), to mine quality single-token and multi-token phrases from the corpus to serve as candidates for taxonomy enrichment. Then, for each class $c$ and each of its parents $p$, we select the top-$k$ terms with the highest affinity scores with $c$ corresponding to $p$, denoted as $T(c,c_{p})$. Then, we aggregate the terms corresponding to each parent of $c$ to get the final enriched class-indicative topical terms for class $c$ as follows
(5) $\small T_{c}=\bigcup_{c_{p}\in Par(c)}T(c,c_{p}).$
3.3. Core Class Refinement with Enriched Taxonomy
With the enriched class-indicative terms for each class, we propose to further utilize them to refine the initial core classes. In this paper, we adopt an embedding-based document-class matching method, which has two distinct advantages in the hierarchical setting. (1) By pre-computing the document and class representations, calculating their similarity is much more efficient than the textual entailment-based method, promoting a more comprehensive score comparison on the entire taxonomy instead of just on a sub-tree. (2) While textual entailment model is trained to predict an entailment score for each pair of document and class, which follows an unknown distribution and is unreliable for corpus-level comparison, the embedding similarity is more general and easier to compare across classes and documents.
Unlike previous methods in flat text classification (Wang et al., 2021a) that use keyword-level embeddings to estimate document and class representations, here, because we already have a set of roughly classified documents for each class, $D^{0}_{c}$, we can directly define their representations based on document-level embeddings. To obtain document representations, we utilize a pretrained Sentence Transformer model (Reimers and Gurevych, 2019) to encode the entire document, which we denote as $\vec{d}$. Then, for each class $c$, we identify a subset of its assigned documents which explicitly mention at least one of the class-indicative keywords and thus most confidently belong to this class, $D_{c}=\{d\in D^{0}_{c}|\exists w\in T_{c},w\in d\}$. Then, we use the average of their document embeddings as the class representation, $\vec{c}=\frac{1}{|D_{c}|}\sum_{d\in D_{c}}\vec{d}$. Finally, we compute the document-class matching score as the cosine similarity between their representations.
Based on the document-class matching scores, we make an observation that the true core classes often have much higher matching scores with the document compared to other classes. Therefore, we use the largest “similarity gap” for each document to identify its core classes. Specifically, for each document $d_{i}\in\mathcal{D}$, we first get a ranked list of classes according to the matching scores, denoted as $[c^{i}_{1},c^{i}_{2},\dots,c^{i}_{|\mathcal{C}|}]$, where $\text{diff}^{\,i}(j):=\cos(\vec{d}_{i},\vec{c}^{i}_{j})-\cos(\vec{d}_{i},\vec{c}^{i}_{j+1})>0$ for $j\in\{1,\dots,|\mathcal{C}|-1\}$. Then, we find the position $m_{i}$ with the highest similarity difference with its next one in the list. After that, we treat the classes ranked above this position as this document’s refined core classes $\mathbb{C}_{i}$, and the corresponding similarity gap as the confidence estimation $conf_{i}$.
(6) $\small\begin{split}conf_{i}=\text{diff}^{\,i}(m_{i}),\quad\mathbb{C}_{i}&=\{c^{i}_{1},c^{i}_{2},\dots,c^{i}_{m_{i}}\},\\ m_{i}=\operatorname*{arg\,max}_{j\in\{1,\dots,|\mathcal{C}|-1\}}&\text{diff}^{\,i}(j).\end{split}$
Finally, we select the set of documents and their refined core classes with the top 50% confidence scores $conf_{i}$, to be the final set of refined core classes for classifier training. Formally, $\mathcal{D}^{\text{core}}=\{d_{i}\}$, where each $d_{i}\in\mathcal{D}^{\text{core}}$ satisfies |$|\{d_{j}|conf_{j}<conf_{i}\}|\geq\frac{1}{2}|\mathcal{D}|$.
3.4. Text Classifier Training with Path-Based Data Augmentation
The final step of TELEClass is to train a hierarchical text classifier using the confident refined core classes. One straightforward way is to directly use the selected core classes as a complete set of pseudo-labeled documents and train a text classifier in a common supervised way. However, such a strategy is ineffective, because the core classes are not comprehensive enough and cannot cover all the classes in the taxonomy. The reason is two-fold. First, the core class selection step is document-centric, meaning that it only focuses on selecting the most accurate core classes for each document, while ignoring whether each class in the taxonomy is ever selected as a core class. Second, the label space in the hierarchical text classification task naturally contains some fine-grained and long-tail classes, so they are often not guaranteed to be selected as core classes due to their low frequency. Empirically, for the two datasets we use in our experiments, Amazon and DBPedia, the missing rates of the selected core classes are 11.6% and 5.4%, respectively. These missing classes will never be used as positive classes in the training process, if we only train the classifier with the selected core classes.
Therefore, to overcome this issue, we further utilize the language generation ability of LLMs to augment the current pseudo-labels. We propose the idea of path-based document generation by LLMs to generate augmented documents for each path from a level-1 node to a leaf node in the taxonomy. By adding the generated documents to the pseudo-labeled data, we can ensure the coverage of the pseudo-data. Moreover, we use a path instead of a single class to guide the LLM generation, because the meaning of lower-level classes is often conditioned on their parents. For example, in Figure 2, a path “hair care” $\rightarrow$ “shampoo” can guide the LLM to generate text about hair shampoo instead of pet shampoo or carpet shampoo that are in different paths.
More specifically, for each path from a level-1 node to a leaf node in the taxonomy, $p=(c_{1},\dots,c_{l})$, we query an LLM to generate a small number of documents (e.g., $q=5$) corresponding to the path $p$ and denote the resulting documents as $d^{p}_{i}$, $i=1,\dots,q$. By generating $q$ documents for each path and treating all the classes in the query path as the pseudo-labels for the generated documents, we get a set of generated data $\mathcal{D}^{\text{gen}}=\{d^{p}_{i}\}_{p\in Path(\mathcal{T}),i=1,\dots,q}$, where $Path(\mathcal{T})$ denotes all such paths in the taxonomy $\mathcal{T}$, and $\mathbb{C}_{p}$ is the set of classes in the path $p$. $\mathcal{D}^{\text{gen}}$ can serve as augmented data to train the text classifier together with the selected refined core classes $\mathcal{D}^{\text{core}}$. Notice that, here we generate pseudo documents for each path instead of just generating for the classes that have no or very few assigned documents. The reason is that, because our core classes are obtained based on static keyword-based features (i.e., the enriched class-indicative topical terms), it may induce certain bias of the matching process into the pseudo-data and affect the learning process (Dong et al., 2023). Therefore, we use an LLM to generate pseudo documents for each path in the taxonomy, which can also serve as regularization of the potential bias.
Now, with two sets of pseudo-data prepared, we are ready to introduce the classifier architecture and the training process.
Classifier architecture. Because we do not focus on proposing a new model for the hierarchical text classification task, we choose to use a standard architecture as our text classifier, including a document encoder, class embeddings, and a matching network.
  • Document encoder: we initialize the document encoder $g(\cdot)$ with a pretrained BERT-base model (Devlin et al., 2019) and use the hidden representation of the [CLS] token from the last layer as the document encoding, $\mathbf{d}_{i}=g(d_{i})$.
  • Class embeddings: we initialize the class embeddings with the class surface name embeddings obtained by a pretrained BERT model (c.f. Section 3.2), denoted by $\mathbf{c}_{j}$. Notice that, here we detach the embeddings from the base model, so only the embeddings will be updated during the training process without back-propagation to the backbone model.
  • Matching Network: we utilize a log-bilinear matching network to calculate the probability of document $d_{i}$ belonging to class $c_{j}$:
    $\small p(c_{j}|d_{i})=\mathcal{P}(y_{j}=1|d_{i})=\sigma(\exp{(\mathbf{c}_{j}^{T}\mathbf{W}\mathbf{d}_{i})}),$
    where $\sigma$ stands for the sigmoid function and $\mathbf{W}$ is a learnable interaction matrix.
Training process. For each document $d_{i}$ tagged with core classes $\mathbb{C}_{i}$, we construct its positive classes $\mathbb{C}^{\text{core}}_{i,+}$ as the union of its core classes and their ancestors in the label taxonomy, and its negative classes $\mathbb{C}^{\text{core}}_{i,-}$ are the ones that are not positive classes or descendants of any core class. This is because the ancestors of confident core classes are also likely to be true labels, and the descendants may not all be negative given that the automatically generated core classes are not optimal. Formally,
$\small\mathbb{C}^{\text{core}}_{i,+}$ $=\mathbb{C}_{i}\cup\left(\cup_{c\in\mathbb{C}_{i}}Anc(c)\right),$
$\mathbb{C}^{\text{core}}_{i,-}$ $=\mathcal{C}-\mathbb{C}^{\text{core}}_{i,+}-\cup_{c\in\mathbb{C}_{i}}Des(c),$
where $Anc(c)$ and $Des(c)$ denote the set of ancestors and descendants and class $c$, respectively. For the LLM-generated documents, we are more confident in its pseudo labels, so we simply treat all the classes in the corresponding path as positive classes and all other classes as negative,
$\small\mathbb{C}^{\text{gen}}_{p,+}=\mathbb{C}_{p},\quad\mathbb{C}^{\text{gen}}_{p,-}=\mathcal{C}-\mathbb{C}_{p}.$
Then, we train a multi-label classifier with the binary cross entropy loss (BCE):
$\mathcal{L}^{\text{core}}=-\sum_{d_{i}\in\mathcal{D}^{\text{core}}}\left(\sum_{c_{j}\in\mathbb{C}^{\text{core}}_{i,+}}\log{p(c_{j}|d_{i})}+\sum_{c_{j}\in\mathbb{C}^{\text{core}}_{i,-}}\log{(1-p(c_{j}|d_{i}))}\right)$
$\mathcal{L}^{\text{gen}}=-\sum_{d^{p}_{i}\in\mathcal{D}^{\text{gen}}}\left(\sum_{c_{j}\in\mathbb{C}^{\text{gen}}_{p,+}}\log{p(c_{j}|d^{p}_{i})}+\sum_{c_{j}\in\mathbb{C}^{\text{gen}}_{p,-}}\log{(1-p(c_{j}|d^{p}_{i}))}\right)$
(7) $\small\mathcal{L}=\mathcal{L}^{\text{core}}+\frac{|\mathcal{D}^{\text{core}}|}{|\mathcal{D}^{\text{gen}}|}\cdot\mathcal{L}^{\text{gen}},$
where $\frac{|\mathcal{D}^{\text{core}}|}{|\mathcal{D}^{\text{gen}}|}$ serves as a scaling factor to balance two sets of pseudo-data and can be exactly calculated as $\frac{|\mathcal{D}|}{2q\cdot|Path(\mathcal{T})|}$.
We train the text-matching model with $\mathcal{L}$ to get the final classifier. Notice that, we do not continue to train the classifier with self-training that is commonly used in previous studies (Meng et al., 2019; Shen et al., 2021), because we want to keep the classifier training part simple. Using self-training may further improve the model performance, which we leave for future exploration. Algorithm 1 summarizes TELEClass.
Input: A corpus $\mathcal{D}$, a label taxonomy $\mathcal{T}$, a textual entailment model $\mathcal{M}$, a sentence encoder $\mathcal{S}$, an LLM $\mathcal{G}$.
Output: A text classifier $F$ that can classify each document into a set of classes in $\mathcal{T}$.
1 // LLM-Enhanced Core Class Annotation;
2 for $d_{i}\in\mathcal{D}$ do
3       $\mathbb{C}^{0}_{i}\leftarrow$ use the LLM $\mathcal{G}$ to select core classes from candidates retrieved by $\mathcal{M}$;
4
5// Corpus-Based Taxonomy Enrichment;
6 for $c\in\mathcal{C}$ do
7       $D^{0}_{c}\leftarrow$ a set of roughly classified documents (Eq. 1);
8       for $c_{p}\in Par(c)$ do
9             $T(c,c_{p})\leftarrow$ top terms ranked by affinity (Eq. 4);
10
11      $T_{c}\leftarrow$ aggregate class-indicative terms with Eq. 5;
12
13// Core Class Refinement with Enriched Taxonomy;
14 $\vec{d}\leftarrow$ document representation $\mathcal{S}(d)$;
15 for $c\in\mathcal{C}$ do
16       $D_{c}\leftarrow$ confident documents by matching $T_{c}$;
17       $\vec{c}\leftarrow$ average document representation in $D_{c}$;
18
19for $d_{i}\in\mathcal{D}$ do
20       $\mathbb{C}_{i},conf_{i}\leftarrow$ refined core classes using $\cos(\vec{d},\vec{c})$ and Eq. 6;
21
[…]
23 // Text Classifier Training with Path-Based Data Augmentation;
24 $\mathcal{D}^{\text{gen}}\leftarrow$ generate $q$ documents for each path using $\mathcal{G}$;
25 $F\leftarrow$ train classifier with $\mathcal{D}^{\text{core}}$, $\mathcal{D}^{\text{gen}}$, and Eq. 7;
26 Return $F$;
Algorithm 1 TELEClass
Table 1. Experiment results on Amazon-531 and DBPedia-298 datasets, evaluated by Example-F1, P@k, and MRR. The best score among zero-shot and weakly-supervised methods is boldfaced. “${}^{\dagger}$” indicates the numbers for these baselines are directly from previous paper (Shen et al., 2021). “—” means the method cannot generate a ranking of predictions and thus MRR cannot be calculated.
Supervision Type Methods Amazon-531 DBPedia-298
Example-F1 P@1 P@3 MRR Example-F1 P@1 P@3 MRR
Zero-Shot ${}^{\dagger}$ 0.4742 0.7144 0.4610 0.6765 0.7871 0.6765
GPT-3.5-turbo 0.5164 0.6807 0.4752 0.4816 0.5328 0.4547
Weakly-Supervised ${}^{\dagger}$ 0.3157 0.5805 0.3115 0.1443 0.2635 0.1443
${}^{\dagger}$ 0.2458 0.5773 0.2517 0.3047 0.5359 0.3048
${}^{\dagger}$ 0.5431 0.7918 0.5414 0.5911 0.7712 0.8621 0.7712 0.8221
${}^{\dagger}$ 0.5934 0.8120 0.5894 0.6332 0.8156 0.8942 0.8156 0.8762
TELEClass 0.6330 0.8439 0.6269 0.6664 0.8684 0.9293 0.8684 0.8880
Fully-Supervised 0.8843 0.9524 0.8758 0.9085 0.9786 0.9945 0.9786 0.9826
4. Experiments
4.1. Experiment Setup
4.1.1. Datasets
We use two public datasets in different domains for evaluation. (1) Amazon-531 (McAuley and Leskovec, 2013) consists of Amazon product reviews and a three-layer label taxonomy of 531 product types, with 29,487 unlabeled training samples and 19,685 testing samples. (2) DBpedia-298 (Lehmann et al., 2015) consists of Wikipedia articles with a three-layer label taxonomy of 298 categories, with 196,665 unlabeled training samples and 49,167 testing samples.
4.1.2. Compared Methods
We compare the following methods on the weakly-supervised hierarchical text classification task.
  • Hier-0Shot-TC (Yin et al., 2019) is a zero-shot approach, which utilizes a pretrained textual entailment model to iterative find the most similar class at each level for a document.
  • GPT-3.5-turbo is a zero-shot approach that queries GPT-3.5-turbo by directly providing all classes in the prompt.
  • Hier-doc2vec (Le and Mikolov, 2014) is a weakly-supervised approach, which first trains document and class representations in the same embedding space, and then iteratively selects the most similar class at each level.
  • WeSHClass (Meng et al., 2019) is a weakly-supervised approach using a set of keywords for each class. It first generates pseudo documents to pretrain text classifiers and then performs self-training.
  • TaxoClass and its variant TaxoClass-NoST (Shen et al., 2021) is a weakly-supervised approach that only uses the class name of each class. It first uses a textual entailment model with a top-down search and corpus-level comparison to select core classes, which are then used as pseudo-training data. We include both its full model, TaxoClass, and its variation TaxoClass-NoST that does not apply self-training on the trained classifier, which is the same as ours.
  • TELEClass is our newly proposed weakly-supervised approach that only uses the class name for each class.
  • Fully-Supervised is a fully-supervised baseline that uses the entire labeled training data to train the text-matching network used in TELEClass.
4.1.3. Evaluation Metrics
Following previous studies (Shen et al., 2021), we utilize the following evaluation metrics: Example-F1, Precision at k (P@k) for $k=1,3$, and Mean Reciprocal Rank (MRR). See Appendix A for definitions of these metrics.
4.1.4. Implementation Details
We use RoBERTa-large-MNLI as the textual entailment model for core class candidate search. We query GPT-3.5-turbo-0613 for both core class annotation and path-based generation. For taxonomy enrichment, we get the term and class name embeddings using a pre-trained BERT-base-uncased. The weight parameter $\alpha$ is set to 0.2, and we select top $k=20$ enriched terms for each class (c.f. Section 3.2). For core class refinement, we get the document and class embeddings using Sentence Transformer all-mpnet-base-v2 (c.f. Section 3.3). We generate $q=5$ documents with path-based generation for each class. The document encoder in the final classifier is initialized with BERT-base-uncased. We train the classifier using AdamW optimizer with a learning rate 5e-5, and the batch size is 64. We include prompts we used for LLM annotation and generation in Appendix B.
4.2. Experimental Results
Table 1 shows the evaluation results of all the compared methods. We make the following observations. (1) Overall, TELEClass achieves the best performance among the compared strong zero-shot and weakly-supervised baselines. (2) By comparing with other weakly-supervised methods, we find that TELEClass significantly outperforms TaxoClass-ST, which is the strongest baseline that also does not use self-training as TELEClass. This further demonstrates the effectiveness of TELEClass. (3) Although LLM (e.g., GPT-3.5-turbo) show superior power in many tasks, naïvely prompting GPT-3.5-turbo in the hierarchical text classification task does not yield promising performance and underperforms strong weakly-supervised text classifier by a large margin. We will conduct a more detailed comparison of LLM prompting for hierarchical text classification in Section 4.4.
Table 2. Performance of TELEClass and its ablations on Amazon-531 and DBPedia-298 datasets. The best score is boldfaced.
Methods Amazon-531 DBPedia-298
Example-F1 P@1 P@3 MRR Example-F1 P@1 P@3 MRR
Gen-Only 0.5151 0.7477 0.5096 0.5357 0.7930 0.9421 0.7930 0.8209
TELEClass-NoEnrich 0.6324 0.8375 0.6263 0.6730 0.6868 0.7726 0.6868 0.7182
TELEClass-NoGen 0.6084 0.8238 0.6024 0.6361 0.8364 0.9146 0.8364 0.8599
TELEClass 0.6330 0.8439 0.6269 0.6664 0.8684 0.9293 0.8684 0.8880
Table 3. Performance comparison of zero-shot LLM prompting. We only report Example-F1 and P@k, because it is not straightforward to get ranking of classes predicted by LLMs for MRR calculation. We also report estimated cost in US dollar (as of the pricing in Feb 2024) and running time in minutes for each method on the entire test set. “${}^{\ddagger}$” indicates that we report the performance on a 1,000-document subset of test data.
Methods Amazon-531 DBPedia-298
Example-F1 P@1 P@3 Est. Cost Est. Time Example-F1 P@1 P@3 Est. Cost Est. Time
GPT-3.5-turbo 0.5164 0.6807 0.4752 \$60 240 mins 0.4816 0.5328 0.4547 \$80 400 mins
GPT-3.5-turbo (level) 0.6621 0.8574 0.6444 \$20 800 mins 0.6649 0.8301 0.6488 \$60 1,000 mins
${}^{\ddagger}$ 0.6994 0.8220 0.6890 \$800 400 mins 0.6054 0.6520 0.5920 \$2,500 1,000 mins
TELEClass 0.6330 0.8439 0.6269 <\$1 3 mins 0.8684 0.9293 0.8684 <\$1 7 mins
4.3. Ablation Studies
To better understand if each component of TELEClass contributes to the entire method, we additionally conduct ablation studies by removing part of TELEClass and reporting the performance. Table 2 shows the results of the following ablations of TELEClass:
  • Gen-Only only uses the augmented documents produced by path-based LLM generation to train the final classifier.
  • TELEClass-NoEnrich does not perform taxonomy enrichment, so directly uses the initial core classes annotated by LLM plus augmented documents to train the final classifier.
  • TELEClass-NoGen does not use the augmented documents produced by path-based LLM generation, so only uses the refined core classes by taxonomy enrichment and embedding-based matching to train the final classifier.
We find that the full model TELEClass achieves the overall best performance among the compared methods, showing the effectiveness of each of its components. Interestingly, we find that taxonomy enrichment and path-based data generation make different levels of contribution on two data sets: path-based data generation brings more improvement on Amazon-531, while taxonomy enrichment contributes more on DBpedia-298. The reasons could be as follows. Amazon-531 has a large label space and more skewed class distribution. We find that around 11.6% of classes are not selected as core classes in Amazon, while this number for DBpedia is 5.4%. This necessitates data augmentation to provide training data that covers each class in the taxonomy. On the other hand, although DBpedia has a smaller label space, its class names are more subtle to distinguish, which can also be demonstrated by the lower performance of zero-shot LLM prompting on DBpedia compared to Amazon-531 (c.f. GPT-3.5-turbo in Table 1). Therefore, by mining class-indicative terms, TELEClass can enrich each class with more clues to facilitate better classification.
4.4. Comparison with Zero-Shot LLM Prompting
In this section, we further compare TELEClass with zero-shot LLM prompting. Because it is not straightforward to get ranked predictions by LLMs, we only report Example-F1 and P@k as performance evaluations. Additionally, we report the estimated cost and time for each method on the entire test set. The cost is reported in US dollar based on the pricing in February 2024. The inference time is reported in minutes, and please be aware that this is just a rough estimation as the actual running time is also dependent on the server condition.
We include the following settings to compare with TELEClass:
  • GPT-3.5-turbo: We include all the classes in the prompt and ask GPT-3.5-turbo model to provide 3 most appropriate classes for a given document.
  • GPT-3.5-turbo (level): We perform level-by-level prompting using GPT-3.5-turbo. Starting from the root node, we ask the model to return one most appropriate class for a given document, and we iteratively prompt the model with the child nodes of the selected node at each level. Notice that this method can only generate a path in the taxonomy, but in the actual multi-label hierarchical classification setting, the true labels may not sit in the same path in the label taxonomy.
  • GPT-4: Similar to the first one, we include all the classes in the prompt and ask GPT-4 to provide 3 most appropriate classes for a given document. Given the limited budget, we only test it on a sampled subset of the test split, which contains 1,000 documents for each dataset.
Table 3 shows the experiment results. We find that TELEClass consistent outperforms all compared methods on DBpedia, while on Amazon, TELEClass underperforms GPT-3.5-turbo (level) and GPT-4 but still being comparable. As for the cost, once trained, TELEClass does not require an additional cost on inference and also has substantially shorter inference time. Prompting LLMs takes longer time and can be prohibitively expensive (e.g., using GPT-4), and the cost will scale up with an increasing size of test data. Also, we find that the level-by-level version of GPT-3.5-turbo consistently outperforms the naïve version, demonstrating the necessity of taxonomy structure. It saves the cost because of the much shorter prompts, but takes longer time due to more queries made per document.
Table 4. Intermediate results on two documents, including selected core classes by different methods, true labels, and the corresponding taxonomy enrichment results. The optimal core class in the true labels is marked with ©.
Dataset Document Core Classes by… True Labels Corr. Enrichment
DBPedia The Lindenhurst Memorial Library (LML) is located in Lindenhurst, New York, and is one of the fifty six libraries that are part of the Suffolk Cooperative Library System … TaxoClass: village TELEClass initial: building TELEClass refined: library library©, agent, educational institution Class: library Top Enrichment: national library, central library, collection, volumes…
Amazon Since mom (89 yrs young) isn’t steady on her feet, we have place these grab bars around the room. It gives her the stability and security she needs. TaxoClass: personal care, health personal care, safety TELEClass initial: daily living aids, medical supplies equipment, safety, TELEClass refined: bathroom aids safety health personal care, medical supplies equipment, bathroom aids safety© Class: bathroom aids safety Top Enrichment: seat, toilet, shower, safety, handles…
4.5. Case studies
To better understand the TELEClass framework, we show some intermediate results of two documents in Table 4, including the core classes selected by (1) the original TaxoClass (Shen et al., 2021) method, (2) TELEClass’s initial core classes selected by LLM, and (3) refined core classes of TELEClass. Besides, we also include the true labels of the documents and the taxonomy enrichment results of the corresponding core class in the table. Overall, we can see that TELEClass’s refined core class is the most accurate. For example, for the first Wikipedia article about a library, TaxoClass selects “village” as the core class, while TELEClass’s initial core class finds a closer one “building” thanks to the power of LLMs. Then, with the enriched classes-indicative features as guidance, TELEClass’s refined core class correctly identifies the optimal core class, which is “library”. In the other example, TELEClass also pinpoints the most accurate core class “bathroom aids safety” while other methods can only find more general or partially relevant classes.
5. Related Work
5.1. Weakly-Supervised Text Classification
Weakly-supervised text classification trains a classifier with limited guidance, aiming to reduce human efforts while maintaining high proficiency. Various sources of weak supervision have been explored, including distant supervision (Gabrilovich and Markovitch, 2007; Chang et al., 2008; Song and Roth, 2014) like knowledge bases, keywords (Agichtein and Gravano, 2000; Tao et al., 2018; Meng et al., 2018, 2020; Wang et al., 2021a; Zhang et al., 2021b), and heuristic rules (Ratner et al., 2016; Badene et al., 2019; Shu et al., 2020). Later, extremely weakly-supervised methods are proposed to solely rely on class label names for generating pseudo-labels and training classifiers. LOTClass (Meng et al., 2020) utilizes MLM-based PLM as a knowledge base for extracting class-indicative keywords. XClass (Wang et al., 2021a) extracts keywords for creating static class representations through clustering. PIEClass (Zhang et al., 2023a) employs PLMs’ zero-shot prompting to obtain pseudo labels with noise-robust iterative ensemble training. MEGClass (Kargupta et al., 2023) acquires contextualized sentence representations to capture topical information at the document level. WOTClass (Wang et al., 2023) focuses on mining and ranking class-indicative keywords from generated classes and extracting classes with overlapping keywords.
5.2. Hierarchical Text Classification
A hierarchy provides a systematic top-to-down structure with inherent semantic relations that can assist in text classification. Typical hierarchical text classification can be categorized into two groups: local approaches and global approaches. Local approaches train multiple classifiers for each node or local structures (Banerjee et al., 2019; Liu et al., 2005; Wehrmann et al., 2018). Global approaches, learn hierarchy structure into a single classifier through recursive regularization (Gopal and Yang, 2013), a graph neural network (GNN)-based encoder (Peng et al., 2018; Huang et al., 2019; Zhou et al., 2020; Shen et al., 2021), or a joint document label embedding space (Chen et al., 2021).
Weak supervision is also studied for hierarchical text classification. WeSHClass (Meng et al., 2019) uses a few keywords or example documents per class and pretrains classifiers with pseudo documents followed by self-training. TaxoClass (Shen et al., 2021) follows the same setting as ours which uses the sole class name of each class as the only supervision. It identifies core classes for each document using a textual entailment model, which is then used to train a multi-label classifier. Additionally, MATCH (Zhang et al., 2021c) and HiMeCat (Zhang et al., 2021a) study how to integrate associated metadata into the label hierarchy for document categorization with weak supervision.
5.3. LLMs as Generators and Annotators
Large language models (LLMs) have demonstrated impressive performance in many downstream tasks and are explored to help low-resource settings by synthesizing data as generators or annotators (Ding et al., 2023). For data generation, few-shot examples (Wang et al., 2021b) or class-conditioned prompts (Meng et al., 2022) are explored for LLM generation and the generated data can be used as pseudo-training data to further fine-tune a small model as the final classifier (Ye et al., 2022). Recently, Yu et al. (2023) propose an attribute-aware topical text classification method that incorporates ChatGPT to generate topic-dependent attributes and topic-independent attributes to reduce topic ambiguity and increase topic diversity for generation. For data annotation, previous works utilize LLMs for unsupervised annotation (Feng et al., 2021), Chain-of-Thought annotation with explanation generation (He et al., 2023), and active annotation (Zhang et al., 2023b).
6. Conclusion and Future Work
In this paper, we propose a new method, TELEClass, for the minimally-supervised hierarchical text classification task with two major contributions. First, we enrich the input label taxonomy with more class-indicative topical terms for each class, which can serve as additional features to understand the classes and facilitate classification. Second, we explore the utilization of LLMs in the hierarchical text classification in two directions: data annotation and data creation. On two public datasets, TELEClass can outperform existing baseline methods substantially, and we further demonstrate its effectiveness through ablation studies. We also conduct a comparative analysis on performance and cost for zero-shot LLM prompting for the hierarchical text classification task. For future works, we plan to generalize TELEClass’s idea into other low-resource text mining tasks with hierarchical label spaces, such as fine-grained entity typing. We also plan to explore how to extend TELEClass into a more complicated label space, such as an extremely large space with millions of labels.
Appendix A Evaluation Metrics
Let $\mathbb{C}^{\text{true}}_{i}$ and $\mathbb{C}^{\text{pred}}_{i}$ denote the set of true labels and the set of predicted labels for document $d_{i}\in\mathcal{D}$, respectively. If a method can generate rankings of classes within $\mathbb{C}^{\text{pred}}_{i}$, we further denote its top-k predicted labels as $\mathbb{C}^{\text{pred}}_{i,k}\subset\mathbb{C}^{\text{pred}}_{i}$. The evaluation metrics are defined as follows:
  • Example-F1 (Sørensen, 1948), which is also called micro-Dice coefficient, evaluates the multi-label classification results without ranking,
[…]
  • Precision at k, or P@k, is a ranking-based metric that evaluates the precision of top-$k$ predicted classes,
[…]
  • Mean Reciprocal Rank, or MRR, is another ranking-based metric, which evaluates the multi-label predictions based on the inverse of true labels’ ranks within predicted classes,
    (10) $\text{MRR}=\frac{1}{|\mathcal{D}|}\sum_{d_{i}\in\mathcal{D}}\frac{1}{|\mathbb{C}_{i}^{\text{true}}|}\sum_{c_{j}\in\mathbb{C}_{i}^{\text{true}}}\frac{1}{\min\{k|c_{j}\in\mathbb{C}^{\text{pred}}_{i,k}\}}.$
Appendix B Prompts for LLM
  • Core class annotation for Amazon-531
    Instruction: You will be provided with an Amazon product review, and your task is to select its product types from the following categories: [Candidate Classes]. Just give the category names as shown in the provided list. Query: [Document]
  • Path-based generation for Amazon-531
    Instruction: Suppose you are an Amazon Reviewer, please generate 5 various and reliable passages following the requirements below: 1. Must generate reviews following the themes of the taxonomy path: [Path]. 2. Must be in length about 100 words. 3. The writing style and format of the text should be a product review. 4. Should keep the generated text to be diverse, specific, and consistent with the given taxonomy path. You should focus on [The Leaf Node on the Path].
  • Core class annotation for DBPedia-298
    Instruction: You will be provided with a Wikipedia article describing an entity at the beginning, and your task is to select its types from the following categories: [Candidate Classes]. Just give the category names as shown in the provided list. Query: [Document]
  • Path-based generation for DBPedia-298
    Instruction: Suppose you are a Wikipedia Contributor, please generate 5 various and reliable passages following the requirements below: 1. Must generate reviews following the themes of the taxonomy path: [Path]. 2. Must be in length about 100 words. 3. The writing style and format of the text should be a Wikipedia page. 4. Should keep the generated text to be diverse, specific, and consistent with the given taxonomy path. You should focus on [The Leaf Node on the Path].

### v2

@@
 TELEClass: Taxonomy Enrichment and LLM-Enhanced Hierarchical Text Classification with Minimal Supervision
-Abstract.
+Abstract
 Hierarchical text classification aims to categorize each document into a set of classes in a label taxonomy.
@@
 On the other hand, previous weakly-supervised hierarchical text classification methods only utilize the raw taxonomy skeleton and ignore the rich information hidden in the text corpus that can serve as additional class-indicative features.
-To tackle the above challenges, we propose TELEClass, Taxonomy Enrichment and LLM-Enhanced weakly-supervised hierarchical text Classification, which (1) automatically enriches the label taxonomy with class-indicative topical terms mined from the corpus to facilitate classifier training and (2) utilizes LLMs for both data annotation and creation tailored for the hierarchical label space.
-Experiments show that TELEClass can outperform previous weakly-supervised hierarchical text classification methods and LLM-based zero-shot prompting methods on two public datasets.
-Weakly-Supervised Text Classification, Hierarchical Text Classification, Taxonomy Enrichment, Large Language Model
-$*{}^{*}$qual Contribution.
-1. Introduction
+To tackle the above challenges, we propose TELEClass, Taxonomy Enrichment and LLM-Enhanced weakly-supervised hierarchical text Classification, which (1) automatically enriches the label taxonomy with class-indicative terms to facilitate classifier training and (2) utilizes LLMs for both data annotation and creation tailored for the hierarchical label space.
+Experiments show that TELEClass can outperform previous weakly-supervised methods and LLM-based zero-shot prompting methods on two public datasets.
+TELEClass: Taxonomy Enrichment and LLM-Enhanced Hierarchical Text Classification with Minimal Supervision
+Yunyi Zhang1, Ruozhen Yangthanks: Equal contribution1, Xueqiang Xu11footnotemark: 11, Rui Li11footnotemark: 12, Jinfeng Xiao1, Jiaming Shen3, Jiawei Han1, 1University of Illinois Urbana-Champaign, 2University of Science and Technology of China, 3Google DeepMind, yzhan238@illinois.edu
+1 Introduction
 Hierarchical text classification, aiming to classify documents into one or multiple classes in a label taxonomy, is an essential and fundamental task in text mining and natural language processing.
-Compared with standard text classification where label space is flat and relatively small (e.g., less than 20 classes), hierarchical text classification is more challenging given the larger and more structured label space and the existence of fine-grained and long-tail classes in the taxonomy.
-Most earlier works tackle this task in fully supervised (You et al., 2019; Huang et al., 2019; Chen et al., 2021) or semi-supervised settings (Berthelot et al., 2019; Gururangan et al., 2019), and different models are proposed to learn from a substantial amount of human-labeled data.
+Compared with standard text classification where label space is flat and relatively small (e.g., less than 20 classes), hierarchical text classification is more challenging given the larger and more structured label space and the existence of fine-grained and long-tail classes.
+Most earlier works tackle this task in fully supervised You et al.
+(2019); Huang et al.
+(2019); Chen et al.
+(2021) or semi-supervised settings Berthelot et al.
+(2019); Gururangan et al.
+(2019), and different models are proposed to learn from a substantial amount of human-labeled data.
 However, acquiring human annotation is often costly, time-consuming, and not scalable.
-Figure 1. An example document tagged with 3 classes. We automatically enrich each node with class-indicative terms and utilize LLMs to facilitate classification.
-Recently, large language models (LLM) such as GPT-4 (OpenAI, 2023) and Llama2 (Touvron et al., 2023) have demonstrated strong performance in flat text classification through zero-shot or few-shot prompting (Sun et al., 2023).
+Figure 1: An example document tagged with 3 classes. We automatically enrich each node with class-indicative terms and utilize LLMs to facilitate classification.
+Recently, large language models (LLM) such as GPT-4 OpenAI (2023) and Claude 3 Anthropic (2024) have demonstrated strong performance in flat text classification through zero-shot or few-shot prompting Sun et al.
+(2023).
 However, applying LLMs in hierarchical settings, with large and structured label spaces, remains challenging.
-Directly including hundreds of classes in prompts is ineffective, leading to structural information loss, increased computational costs, and diminished clarity for LLMs in focusing on critical information.
+Directly including hundreds of classes in prompts is ineffective, leading to structural information loss, increased computational costs, and diminished clarity for LLMs at distinguishing class-specific information.
 Along another line of research, Meng et al.
@@
 (2021) study the hierarchical text classification with minimal supervision, which takes the class name as the only supervision signal.
-Specifically, they introduce TaxoClass, a method that generates pseudo-labels with a textual entailment model for classifier training.
-However, this method overlooks additional class-relevant features in the text and suffers from unreliable pseudo-label selection due to the entailment model’s domain shift.
-In this study, we advance minimally supervised hierarchical text classification by integrating corpus-based taxonomy enrichment and leveraging LLMs tailored for the hierarchical label structure.
-First, we enrich the existing label taxonomy skeleton with automatically extracted class-specific topical terms from the corpus, enhancing supervision signals and thus improving the pseudo-label quality for classifier training.
-Second, we explore the application of LLMs in hierarchical text classification from two perspectives.
-Specifically, we enhance LLM annotation efficiency and effectiveness through a taxonomy-guided candidate search and optimize LLM-generated document accuracy by using taxonomy paths to create more precise pseudo-data.
+Specifically, they introduce TaxoClass which generates pseudo labels with a textual entailment model for classifier training.
+However, this method overlooks additional class-relevant features in the text and suffers from the unreliable pseudo label selection due to domain shift of the entailment model.
+In this study, we advance minimally supervised hierarchical text classification by enriching the weak supervision with additional textual clues and leveraging LLMs tailored for the hierarchical label structure.
+First, we enrich the existing label taxonomy with class-specific terms derived from two sources: LLM generation and automated extraction from the corpus.
+For example, the “conditioner” class in Figure 1 is enriched with key terms like “moisture” and “soft hair”, which distinguish it from other classes.
+These terms enhance the supervision signal by combining knowledge from LLMs and text corpus and improve the pseudo label quality for classifier training.
+Second, we further explore the application of LLMs in hierarchical text classification from two perspectives.
+Specifically, we enhance LLM annotation efficiency and effectiveness through a taxonomy-guided candidate search and optimize LLM document generation to create more precise pseudo data by using taxonomy paths.
 Leveraging the above ideas, we introduce TELEClass: Taxonomy Enrichment and LLM-Enhanced weakly-supervised hierarchical text Classification.
-TELEClass consists of four major steps: (1) LLM-Enhanced Core Class Annotation, where we identify document “core classes” (i.e., fine-grained classes that most accurately describe the documents) by first using a textual entailment model for top-down candidate search on the taxonomy and then prompting an LLM to select the most precise core classes.
-(2) Corpus-Based Taxonomy Enrichment, where we analyze the taxonomy structure and the roughly classified documents to identify class-indicative topical terms through semantic and statistical analysis, enriching the label taxonomy with more features.
-For example, the “conditioner” class in Figure 1 is enriched with key terms like “moisture” and “soft hair”, which distinguish it from its sibling classes.
+TELEClass consists of four major steps: (1) LLM-Enhanced Core Class Annotation, where we identify document “core classes” (i.e., fine-grained classes that most accurately describe the documents) by first enriching the taxonomy with LLM-generated key terms and then finding candidate classes with a top-down tree search algorithm for LLM to select the most precise core classes.
+(2) Corpus-Based Taxonomy Enrichment, where we analyze the taxonomy structure to additionally identify class-indicative topical terms through semantic and statistical analysis on the corpus.
 (3) Core Class Refinement with Enriched Taxonomy, where we embed documents and classes based on the enriched label taxonomy and refine the initially selected core classes by identifying the most similar classes for each document.
 (4) Text Classifier Training with Path-Based Data Augmentation, where we sample label paths from the taxonomy and guide the LLM to generate pseudo documents most accurately describing these fine-grained classes.
-Finally, we train the text classifier on two types of pseudo-labels, the core classes and the generated data, with a simple text-matching network and multi-label training strategy.
+Finally, we train the text classifier on two types of pseudo labels, the core classes and the generated data, with a simple text matching network and multi-label training strategy.
 The contributions of this paper are summarized as follows: (1) We propose TELEClass, a new method for minimally supervised hierarchical text classification, which requires only the class names of the label taxonomy as supervision to train a multi-label text classifier.
-(2) We propose to enrich the label taxonomy with class-indicative topical terms mined from the text corpus, based on which we utilize an embedding-based document-class matching method to improve the pseudo-label quality.
-(3) We study two ways of adopting large language models to the hierarchical text classification setting, which can improve the pseudo-label quality and solve the data scarcity issue for fine-grained classes.
+(2) We propose to enrich the label taxonomy with class-indicative terms, based on which we utilize an embedding-based document-class matching method to improve the pseudo label quality.
+(3) We study two ways of adopting large language models to help hierarchical text classifier training, which can improve the pseudo label quality and solve the data scarcity issue for fine-grained classes.
 (4) Extensive experiments on two datasets show that TELEClass can outperform strong weakly-supervised hierarchical text classification methods and zero-shot LLM prompting methods.
-[^1] Code and datasets are available upon request now and will later be available online.
-2. Problem Definition
-Following Shen et al.
-(2021), we define the minimally-supervised hierarchical text classification task as to train a text classifier that can categorize each document into multiple nodes on a label taxonomy by using the name of each node as the only supervision.
+2 Problem Definition
+The minimally-supervised hierarchical text classification task aims to train a text classifier that can categorize each document into multiple nodes on a label taxonomy by using the name of each node as the only supervision Shen et al.
+(2021).
 For example, in Figure 1, the input document is classified as “hair care”, “shampoo”, and “scalp treatment”.
 […]
-3. Methodology
+Figure 2: Overview of the TELEClass framework.
+We assume the label taxonomy to be a DAG instead of a tree, because it aligns better with real applications as one node can have multiple parents with different meanings.
+It is also more challenging because the classifier needs to assign a document multiple labels in different levels and paths.
+3 Methodology
 In this section, we will introduce TELEClass consisting of the following modules: (1) LLM-enhanced core class annotation, (2) corpus-based taxonomy enrichment, (3) core class refinement with enriched taxonomy, and (4) text classifier training with path-based data augmentation.
 Figure 2 shows an overview of TELEClass.
-Figure 2. Overview of the TELEClass framework.
-3.1. LLM-enhanced Core Class Annotation
-Inspired by previous studies, we first tag each document with its “core classes”, which are defined as a set of classes that can describe the document most accurately (Shen et al., 2021).
-This process also mimics the process of human performing hierarchical text classification: first, select a set of most essential classes for the document and then trace back to their ancestors to complete the labeling.
+3.1 LLM-Enhanced Core Class Annotation
+Inspired by previous studies, we first tag each document with its “core classes”, which are defined as a set of classes that can describe the document most accurately Shen et al.
+(2021).
+This process also mimics the process of human performing hierarchical text classification: first select a set of most essential classes for the document and then trace back to their relevant classes to complete the labeling.
 For example, in Figure 1, by first tagging the document with “shampoo” and “scalp treatment”, we can easily find its complete set of classes.
-To mine the core classes for each document, existing methods utilize a textual entailment model (e.g., RoBERTa-large-MNLI) and combine tree-based top-down candidate searching with corpus-level statistical analysis for core class selection.
-Specifically, the similarity score between a document and a class is defined as the entailment score from the document to a class-specific hypothesis such as “This example is [Class Name]”.
-These similarity scores facilitate a top-down search in the taxonomy to shortlist candidate classes, which are then evaluated across the corpus to determine the most reliable core classes.
-However, there are two limitations of this method.
-First, the textual entailment model, not being directly trained for document classification, offers inconsistent zero-shot performance, particularly in distinguishing closely related, fine-grained classes within a hierarchical structure.
-Second, given a pair of premise (document) and hypothesis, the entailment model is trained to predict a distribution over three classes, namely entailment, neutral, and contradiction.
-Therefore, it is questionable whether the generated scores can support corpus-level comparison, because the scores between different pairs may not be directly comparable.
-Both limitations will lead to noisy core class selection and adversely impact the classifier performance.
-In this work, we propose to utilize an LLM to enhance the core class annotation step.
-These LLMs are trained on very large textual data and human preference data and shown to have superior zero-shot ability for the document classification task with a small and flat label space (Sun et al., 2023).
-However, LLMs cannot effectively understand the hierarchical label space, which is large and structured and cannot be easily represented in a prompt.
-Therefore, we precede the LLM annotation with the top-down candidate search method using a textual entailment model.
-Specifically, for each document $d_{i}$, we start from the Root node at level 0 of the taxonomy by adding it into a queue.
-For each class at level $l$ in the queue, we select $l+3$ of its children that have the highest similarity to document $d_{i}$, where the similarity score $sim^{\text{entail}}(c,d_{i})$ is obtained from a textual entailment model.
-Then, among all the selected children at level $l+1$, we keep the top $(l+2)^{2}$ of them with the highest path-based similarity scores and add to the queue, where the path-based similarity $sim^{\text{path}}(c,d_{i})$ is defined recursively as,
-$sim^{\text{path}}(\text{Root},d)$ $=1,$
-$sim^{\text{path}}(c,d)$ $=\max_{c_{p}\in Par(c)}sim^{\text{path}}(c_{p},d_{i})\cdot sim^{\text{entail}}(c,d_{i}),$
-where $Par(c)$ is the set of all parents of $c$.
-This process finishes when there is no class in the queue, and all the nodes (other than Root) ever pushed into the queue are treated as candidate core classes for document $d_{i}$.
-To use an LLM to annotate core classes, we first construct an instruction for each document, which asks the LLM to select the most accurate classes for the document from its candidate core classes.
-Then, the document is fed to the LLM as a query and it will generate a list of core classes according to the document and candidate core classes.
-Our LLM-enhanced core class annotation step will assign an initial set of core classes (denoted as $\mathbb{C}^{0}_{i}$) for each document $d_{i}\in\mathcal{D}$.
-In the next two steps, we introduce how TELEClass leverages the initial core classes to enrich the label taxonomy with class-indicative features.
-This enriched taxonomy is subsequently employed to refine the initial core classes, utilizing embedding-based similarity scores that are more comparable across documents.
-3.2. Corpus-Based Taxonomy Enrichment
-Due to the hierarchical nature of taxonomy, fined-grained classes at lower levels are often hard to distinguish with only class label names.
-Therefore, we propose to enrich the weak supervision, i.e.
-the label taxonomy, with a set of class-indicative topical terms for each class.
-These enriched terms can help better interpret the class and thus facilitate text classification.
-For example, in Figure 2, “shampoo” and “conditioner” are two fine-grained classes that are similar to each other.
+In this work, we propose to enhance the core class annotation process of previous methods with the power of LLMs.
+To utilize LLMs for core class annotation, we apply a structure-aware candidate core class selection method to reduce the label space for each document.
+This step is necessary because LLMs can hardly comprehend a large, structured hierarchical label space that cannot be easily represented in a prompt.
+We first define a similarity score between a class and a document that will be used in candidate selection.
+To better capture the semantics, we propose to use LLMs to generate a set of class relevant keywords to enrich the raw taxonomy structure and consolidate the meaning of each class.
+For example, “shampoo” and “conditioner” are two fine-grained classes that are similar to each other.
 We can effectively separate the two classes by identifying a set of class-specific terms such as “flakes” for “shampoo” and “moisture” for “conditioner”.
-In this work, we combine statistical document occurrence features and semantic similarity and further contrast a class with its siblings to identify a set of key terms that can distinguish it from its most similar nodes in the taxonomy.
-Notice that, because the input taxonomy $\mathcal{T}$ is a directed acyclic graph rather than a simple tree, a node $c$ can have multiple sets of siblings, each of which corresponds to one of $c$’s parent nodes.
-Besides, because we have already tagged each document with its core classes, we can use them as rough classification results and form a document cluster for each class to help the enrichment step.
-Formally, given a class $c\in\mathcal{C}$ and its siblings corresponding to one of its parents $c_{p}$, $Sib(c,c_{p})=\{c^{\prime}\in\mathcal{C}|\langle c_{p},c^{\prime}\rangle\in\mathcal{R}\}$, $\mathbf{c}_{p}\in Par(c)$, we find a set of class-indicative topical terms of $c$ corresponding to $c_{p}$, denoted as $T(c,c_{p})=\{t_{1},t_{2},\dots,t_{k}\}$.
+We prompt an LLM to enrich the raw label taxonomy with a set of key terms for each class, denoted as $T^{\text{LLM}}_{c}$ for the class $c$.
+To ensure the generated terms can uniquely identify $c$, we ask the LLM to generate terms that are relevant to $c$ and its parent while irrelevant to siblings of $c$.
+With this enriched set of terms for each class, we define the similarity score between a document $d$ and a class $c$ as the maximum cosine similarity with the key terms:
+$sim(c,d)=\max_{t\in T^{\text{LLM}}_{c}}\cos(\vec{t},\vec{d}),\vspace{-0.5em}$ (1)
+where $\vec{\circ}$ denotes a vector representation by a pre-trained semantic encoder (e.g., Sentence Transformer Reimers and Gurevych (2019)).
+This newly defined similarity measure is then used for candidate core class selection.
+Given a document, we start from the root node at level $l=0$, select the $l+3$ most similar children classes to the document at level $l$ using the similarity score defined above, and continue to the next level with only the selected classes.
+The increasing number of selected nodes accounts for the growing number of classes when going deeper into the taxonomy.
+Finally, all the classes ever selected in this process will be the candidate core classes for this document, which share the most similarity with the document according to the label hierarchy.
+[^1] Refer to Shen et al.
+(2021) for more details on the tree search algorithm.
+Finally, we instruct an LLM to select the core classes for each document from the selected candidates, which produces an initial set of core classes (denoted as $\mathbb{C}^{0}_{i}$) for each document $d_{i}\in\mathcal{D}$.
+3.2 Corpus-Based Taxonomy Enrichment
+In the previous step, we enrich the raw taxonomy structure with LLM-generated key terms, which are derived from the general knowledge of LLMs but may not accurately reflect the corpus-specific knowledge.
+Therefore, we propose further enriching the classes with class-indicative terms mined from the text corpus.
+By doing this, we can combine the general knowledge of LLMs and corpus-specific knowledge to better enhance the very weak supervision, which is essential for correctly understanding fine-grained classes that are hard to distinguish.
+Formally, given a class $c\in\mathcal{C}$ and its siblings corresponding to one of its parents $c_{p}$, $Sib(c,c_{p})=\{c^{\prime}\in\mathcal{C}|\langle c_{p},c^{\prime}\rangle\in\mathcal{R}\}$, $\mathbf{c}_{p}\in Par(c)$, we find a set of corpus-based class-indicative terms of $c$ corresponding to $c_{p}$, denoted as $T(c,c_{p})=\{t_{1},t_{2},\dots,t_{k}\}$.
 Each term in $T(c,c_{p})$ can signify the class $c$ and distinguish it from its siblings under $c_{p}$.
-To identify these class-indicative terms, we construct a set of documents $D^{0}_{c}$ for each class $c$, containing all the documents whose initial core classes contain $c$ or its descendants,
-(1) $\small D^{0}_{c}=\left\{d_{i}\in D|\exists c^{\prime}\in\mathbb{C}^{0}_{i},c^{\prime}\in\{c\}\cup Des(c)\right\},$
-where $Des(c)$ is the set of all descedants of $c$. We consider the following 3 factors for class-indicative term selection: popularity, distinctiveness, and semantic similarity.
-Popularity.
-First, a class-indicative term of $c$ should be frequently mentioned by the documents about $c$, denoted as its popularity within $c$.
-We calculate the popularity of a term $t$ within a class $c$ as,
-(2) $\small pop(t,c)=\log(1+df(t,D^{0}_{c})),$
-where the document frequency $df(t,D)$ stands for the number of documents in $D$ that mention $t$.
-Distinctiveness.
-Second, a class-indicative term $t$ for a class $c$ should also be infrequent in its siblings so that a document containing $t$ likely indicates that it is about $c$ instead of $c$’s siblings.
-We define the distinctiveness score of a term $t$ for a class $c$, corresponding to its siblings under $c_{p}$, as,
-(3) $\small dist(t,c,c_{p})=\frac{\exp(BM25(t,D^{0}_{c}))}{1+\sum_{c^{\prime}\in Sib(c,c_{p})}\exp(BM25(t,D^{0}_{c^{\prime}}))},$
-where $BM25(\cdot,\cdot)$ denotes the BM25 relevance function (Robertson and Zaragoza, 2009).
-[…]
-Finally, we define the affinity score between a term $t$ and a class $c$ corresponding to parent $p$ to be the geometric mean of the above scores, with a weight parameter $\alpha$,
-(4) $\small aff(t,c,c_{p})=pop(t,c)^{\alpha}\cdot dist(t,c,c_{p})^{1-\alpha}\cdot\cos(\mathbf{t},\mathbf{c}).$
-Taxonomy Enrichment.
-We first apply a phrase mining tool, AutoPhrase (Shang et al., 2018), to mine quality single-token and multi-token phrases from the corpus to serve as candidates for taxonomy enrichment.
-Then, for each class $c$ and each of its parents $p$, we select the top-$k$ terms with the highest affinity scores with $c$ corresponding to $p$, denoted as $T(c,c_{p})$.
-Then, we aggregate the terms corresponding to each parent of $c$ to get the final enriched class-indicative topical terms for class $c$ as follows
-(5) $\small T_{c}=\bigcup_{c_{p}\in Par(c)}T(c,c_{p}).$
-3.3. Core Class Refinement with Enriched Taxonomy
+We first collect a set of relevant documents $D^{0}_{c}\subset\mathcal{D}$ for each class $c$, which contains all the documents whose initial core classes contain $c$ or its descendants.
+Then, inspired by Tao et al.
+(2018); Zhang et al.
+(2023b), we consider the following three factors for class-indicative term selection: (1) Popularity: a class-indicative term $t$ of a class $c$ should be frequently mentioned by its relevant documents, which is quantified by the document frequency $\log(1+df(t,D^{0}_{c}))$.
+(2) Distinctiveness: a class-indicative term $t$ for a class $c$ should be infrequent in its siblings, which is quantified as the softmax of $BM25(t,D^{0}_{c})$ relevance scores over the set of siblings, $Sib(c,c_{p})$.
+(3) Semantic similarity: a class-indicative term should also be semantically similar to the class name, which is quantified as the cosine similarity between their embeddings derived from a pre-trained encoder (e.g., BERT Devlin et al.
+(2019)).
+Finally, we define the affinity score between a term $t$ and a class $c$ corresponding to parent $p$ to be the geometric mean of the above scores, denoted as $\text{aff}(t,c,c_{p})$.
+To enrich the taxonomy, we first apply a phrase mining tool, AutoPhrase Shang et al.
+(2018), to mine quality single-token and multi-token phrases from the corpus as candidate terms [^2] Our method is flexible with any kinds of phrase mining methods like Gu et al.
+(2021)..
+Then, for each class $c$ and each of its parents $c_{p}$, we select the top-$k$ terms with the highest affinity scores with $c$ corresponding to $p$, denoted as $T(c,c_{p})$.
+Then, we take the union of these corpus-based terms together with the LLM-generated terms in the previous step to get the final enriched class-indicative terms for class $c$, $T_{c}=\left(\bigcup_{c_{p}\in Par(c)}T(c,c_{p})\right)\bigcup T^{\text{LLM}}_{c}$.
+3.3 Core Class Refinement with Enriched Taxonomy
 With the enriched class-indicative terms for each class, we propose to further utilize them to refine the initial core classes.
-In this paper, we adopt an embedding-based document-class matching method, which has two distinct advantages in the hierarchical setting.
-(1) By pre-computing the document and class representations, calculating their similarity is much more efficient than the textual entailment-based method, promoting a more comprehensive score comparison on the entire taxonomy instead of just on a sub-tree.
-(2) While textual entailment model is trained to predict an entailment score for each pair of document and class, which follows an unknown distribution and is unreliable for corpus-level comparison, the embedding similarity is more general and easier to compare across classes and documents.
-Unlike previous methods in flat text classification (Wang et al., 2021a) that use keyword-level embeddings to estimate document and class representations, here, because we already have a set of roughly classified documents for each class, $D^{0}_{c}$, we can directly define their representations based on document-level embeddings.
-To obtain document representations, we utilize a pretrained Sentence Transformer model (Reimers and Gurevych, 2019) to encode the entire document, which we denote as $\vec{d}$.
-Then, for each class $c$, we identify a subset of its assigned documents which explicitly mention at least one of the class-indicative keywords and thus most confidently belong to this class, $D_{c}=\{d\in D^{0}_{c}|\exists w\in T_{c},w\in d\}$.
+In this paper, we adopt an embedding-based document-class matching method.
+Unlike previous methods in flat text classification Wang et al.
+(2021a) that use keyword-level embeddings to estimate document and class representations, here, we are able to define class representations directly based on document-level embeddings thanks to the rough class assignments we created in the core class annotation step (c.f.
+Sec.
+3.1).
+To obtain document representations, we utilize a pre-trained Sentence Transformer model Reimers and Gurevych (2019) to encode the entire document, which we denote as $\vec{d}$.
+Then, for each class $c$, we identify a subset of its assigned documents that explicitly mention at least one of the class-indicative keywords and thus most confidently belong to this class, $D_{c}=\{d\in D^{0}_{c}|\exists w\in T_{c},w\in d\}$.
 Then, we use the average of their document embeddings as the class representation, $\vec{c}=\frac{1}{|D_{c}|}\sum_{d\in D_{c}}\vec{d}$.
@@
 After that, we treat the classes ranked above this position as this document’s refined core classes $\mathbb{C}_{i}$, and the corresponding similarity gap as the confidence estimation $conf_{i}$.
-(6) $\small\begin{split}conf_{i}=\text{diff}^{\,i}(m_{i}),\quad\mathbb{C}_{i}&=\{c^{i}_{1},c^{i}_{2},\dots,c^{i}_{m_{i}}\},\\ m_{i}=\operatorname*{arg\,max}_{j\in\{1,\dots,|\mathcal{C}|-1\}}&\text{diff}^{\,i}(j).\end{split}$
-Finally, we select the set of documents and their refined core classes with the top 50% confidence scores $conf_{i}$, to be the final set of refined core classes for classifier training.
-Formally, $\mathcal{D}^{\text{core}}=\{d_{i}\}$, where each $d_{i}\in\mathcal{D}^{\text{core}}$ satisfies |$|\{d_{j}|conf_{j}<conf_{i}\}|\geq\frac{1}{2}|\mathcal{D}|$.
-3.4. Text Classifier Training with Path-Based Data Augmentation
+$\begin{split}conf_{i}=\text{diff}^{\,i}(m_{i}),\quad\mathbb{C}_{i}&=\{c^{i}_{1},\dots,c^{i}_{m_{i}}\},\\ m_{i}=\operatorname*{arg\,max}_{j\in\{1,\dots,|\mathcal{C}|-1\}}&\text{diff}^{\,i}(j).\end{split}$ (2)
+Finally, we select the top 75% of documents $d_{i}$ and their refined core classes with the highest confidence scores $conf_{i}$, denoted as $\mathcal{D}^{\text{core}}$.
+3.4 Text Classifier Training with Path-Based Data Augmentation
 The final step of TELEClass is to train a hierarchical text classifier using the confident refined core classes.
@@
 However, such a strategy is ineffective, because the core classes are not comprehensive enough and cannot cover all the classes in the taxonomy.
-The reason is two-fold.
-First, the core class selection step is document-centric, meaning that it only focuses on selecting the most accurate core classes for each document, while ignoring whether each class in the taxonomy is ever selected as a core class.
-Second, the label space in the hierarchical text classification task naturally contains some fine-grained and long-tail classes, so they are often not guaranteed to be selected as core classes due to their low frequency.
-Empirically, for the two datasets we use in our experiments, Amazon and DBPedia, the missing rates of the selected core classes are 11.6% and 5.4%, respectively.
-These missing classes will never be used as positive classes in the training process, if we only train the classifier with the selected core classes.
-Therefore, to overcome this issue, we further utilize the language generation ability of LLMs to augment the current pseudo-labels.
-We propose the idea of path-based document generation by LLMs to generate augmented documents for each path from a level-1 node to a leaf node in the taxonomy.
-By adding the generated documents to the pseudo-labeled data, we can ensure the coverage of the pseudo-data.
+This is because the hierarchical label space naturally contains fine-grained and long-tail classes, and they are often not guaranteed to be selected as core classes due to their low frequency.
+Empirically, for the two datasets we use in our experiments, Amazon and DBPedia, the percentages of classes never selected as core classes are 11.6% and 5.4%, respectively.
+These missing classes will never be used as positive classes in the training process if we only train the classifier with the selected core classes.
+Therefore, to overcome this issue, we propose the idea of path-based document generation by LLMs to generate a small number of augmented documents (e.g., $q=5$) for each distinct path from a level-1 node to a leaf node in the taxonomy.
+By adding the generated documents to the pseudo-labeled data, we can ensure that each class of the taxonomy will be a positive class of at least $q$ documents.
 Moreover, we use a path instead of a single class to guide the LLM generation, because the meaning of lower-level classes is often conditioned on their parents.
 For example, in Figure 2, a path “hair care” $\rightarrow$ “shampoo” can guide the LLM to generate text about hair shampoo instead of pet shampoo or carpet shampoo that are in different paths.
-More specifically, for each path from a level-1 node to a leaf node in the taxonomy, $p=(c_{1},\dots,c_{l})$, we query an LLM to generate a small number of documents (e.g., $q=5$) corresponding to the path $p$ and denote the resulting documents as $d^{p}_{i}$, $i=1,\dots,q$.
-By generating $q$ documents for each path and treating all the classes in the query path as the pseudo-labels for the generated documents, we get a set of generated data $\mathcal{D}^{\text{gen}}=\{d^{p}_{i}\}_{p\in Path(\mathcal{T}),i=1,\dots,q}$, where $Path(\mathcal{T})$ denotes all such paths in the taxonomy $\mathcal{T}$, and $\mathbb{C}_{p}$ is the set of classes in the path $p$.
-$\mathcal{D}^{\text{gen}}$ can serve as augmented data to train the text classifier together with the selected refined core classes $\mathcal{D}^{\text{core}}$.
-Notice that, here we generate pseudo documents for each path instead of just generating for the classes that have no or very few assigned documents.
-The reason is that, because our core classes are obtained based on static keyword-based features (i.e., the enriched class-indicative topical terms), it may induce certain bias of the matching process into the pseudo-data and affect the learning process (Dong et al., 2023).
-Therefore, we use an LLM to generate pseudo documents for each path in the taxonomy, which can also serve as regularization of the potential bias.
-Now, with two sets of pseudo-data prepared, we are ready to introduce the classifier architecture and the training process.
+To promote data diversity, we make one LLM query for each path and ask it to generate $q$ diverse documents.
+Appx.
+C shows the prompts we used.
+We denote the generated documents as $\mathcal{D}^{\text{gen}}$.
+Now, with two sets of data, the pseudo-labeled documents $\mathcal{D}^{\text{core}}$ and LLM-generated documents $\mathcal{D}^{\text{gen}}$, we are ready to introduce the classifier architecture and the training process.
 Classifier architecture.
-Because we do not focus on proposing a new model for the hierarchical text classification task, we choose to use a standard architecture as our text classifier, including a document encoder, class embeddings, and a matching network.
-  • Document encoder: we initialize the document encoder $g(\cdot)$ with a pretrained BERT-base model (Devlin et al., 2019) and use the hidden representation of the [CLS] token from the last layer as the document encoding, $\mathbf{d}_{i}=g(d_{i})$.
-  • Class embeddings: we initialize the class embeddings with the class surface name embeddings obtained by a pretrained BERT model (c.f.
-Section 3.2), denoted by $\mathbf{c}_{j}$.
-Notice that, here we detach the embeddings from the base model, so only the embeddings will be updated during the training process without back-propagation to the backbone model.
-  • Matching Network: we utilize a log-bilinear matching network to calculate the probability of document $d_{i}$ belonging to class $c_{j}$:
-    $\small p(c_{j}|d_{i})=\mathcal{P}(y_{j}=1|d_{i})=\sigma(\exp{(\mathbf{c}_{j}^{T}\mathbf{W}\mathbf{d}_{i})}),$
-    where $\sigma$ stands for the sigmoid function and $\mathbf{W}$ is a learnable interaction matrix.
+We use a simple text matching network Shen et al.
+(2021) as our model architecture, which includes a document encoder initialized with a pre-trained BERT-base model Devlin et al.
+(2019) and a log-bilinear matching network.
+Class representations are initialized by class name embeddings (c.f.
+Sec.
+3.2).
 Training process.
-For each document $d_{i}$ tagged with core classes $\mathbb{C}_{i}$, we construct its positive classes $\mathbb{C}^{\text{core}}_{i,+}$ as the union of its core classes and their ancestors in the label taxonomy, and its negative classes $\mathbb{C}^{\text{core}}_{i,-}$ are the ones that are not positive classes or descendants of any core class.
+For each document with refined core classes, we construct its positive classes as the union of its refined core classes and their ancestors in the label taxonomy, and its negative classes are the ones that are not positive classes or descendants of any core class.
 This is because the ancestors of confident core classes are also likely to be true labels, and the descendants may not all be negative given that the automatically generated core classes are not optimal.
-Formally,
-$\small\mathbb{C}^{\text{core}}_{i,+}$ $=\mathbb{C}_{i}\cup\left(\cup_{c\in\mathbb{C}_{i}}Anc(c)\right),$
-$\mathbb{C}^{\text{core}}_{i,-}$ $=\mathcal{C}-\mathbb{C}^{\text{core}}_{i,+}-\cup_{c\in\mathbb{C}_{i}}Des(c),$
-where $Anc(c)$ and $Des(c)$ denote the set of ancestors and descendants and class $c$, respectively.
-For the LLM-generated documents, we are more confident in its pseudo labels, so we simply treat all the classes in the corresponding path as positive classes and all other classes as negative,
-$\small\mathbb{C}^{\text{gen}}_{p,+}=\mathbb{C}_{p},\quad\mathbb{C}^{\text{gen}}_{p,-}=\mathcal{C}-\mathbb{C}_{p}.$
-Then, we train a multi-label classifier with the binary cross entropy loss (BCE):
-$\mathcal{L}^{\text{core}}=-\sum_{d_{i}\in\mathcal{D}^{\text{core}}}\left(\sum_{c_{j}\in\mathbb{C}^{\text{core}}_{i,+}}\log{p(c_{j}|d_{i})}+\sum_{c_{j}\in\mathbb{C}^{\text{core}}_{i,-}}\log{(1-p(c_{j}|d_{i}))}\right)$
-$\mathcal{L}^{\text{gen}}=-\sum_{d^{p}_{i}\in\mathcal{D}^{\text{gen}}}\left(\sum_{c_{j}\in\mathbb{C}^{\text{gen}}_{p,+}}\log{p(c_{j}|d^{p}_{i})}+\sum_{c_{j}\in\mathbb{C}^{\text{gen}}_{p,-}}\log{(1-p(c_{j}|d^{p}_{i}))}\right)$
-(7) $\small\mathcal{L}=\mathcal{L}^{\text{core}}+\frac{|\mathcal{D}^{\text{core}}|}{|\mathcal{D}^{\text{gen}}|}\cdot\mathcal{L}^{\text{gen}},$
-where $\frac{|\mathcal{D}^{\text{core}}|}{|\mathcal{D}^{\text{gen}}|}$ serves as a scaling factor to balance two sets of pseudo-data and can be exactly calculated as $\frac{|\mathcal{D}|}{2q\cdot|Path(\mathcal{T})|}$.
-We train the text-matching model with $\mathcal{L}$ to get the final classifier.
-Notice that, we do not continue to train the classifier with self-training that is commonly used in previous studies (Meng et al., 2019; Shen et al., 2021), because we want to keep the classifier training part simple.
+For the LLM-generated documents, we are more confident in their labels, so we simply treat all the classes in the corresponding label path as positive classes and all other classes as negative.
+We train the text-matching model with the standard binary cross-entropy loss.
+The loss terms of two sets of data are weighted by their relative size, $\frac{|\mathcal{D}^{\text{core}}|}{|\mathcal{D}^{\text{gen}}|}$.
+Notice that we do not continue training the classifier with self-training that is commonly used in previous studies Meng et al.
+(2019); Shen et al.
+(2021).
 Using self-training may further improve the model performance, which we leave for future exploration.
-Algorithm 1 summarizes TELEClass.
-Input: A corpus $\mathcal{D}$, a label taxonomy $\mathcal{T}$, a textual entailment model $\mathcal{M}$, a sentence encoder $\mathcal{S}$, an LLM $\mathcal{G}$.
-Output: A text classifier $F$ that can classify each document into a set of classes in $\mathcal{T}$.
-1 // LLM-Enhanced Core Class Annotation;
-2 for $d_{i}\in\mathcal{D}$ do
-3       $\mathbb{C}^{0}_{i}\leftarrow$ use the LLM $\mathcal{G}$ to select core classes from candidates retrieved by $\mathcal{M}$;
-4
-5// Corpus-Based Taxonomy Enrichment;
-6 for $c\in\mathcal{C}$ do
-7       $D^{0}_{c}\leftarrow$ a set of roughly classified documents (Eq. 1);
-8       for $c_{p}\in Par(c)$ do
-9             $T(c,c_{p})\leftarrow$ top terms ranked by affinity (Eq. 4);
-10
-11      $T_{c}\leftarrow$ aggregate class-indicative terms with Eq. 5;
-12
-13// Core Class Refinement with Enriched Taxonomy;
-14 $\vec{d}\leftarrow$ document representation $\mathcal{S}(d)$;
-15 for $c\in\mathcal{C}$ do
-16       $D_{c}\leftarrow$ confident documents by matching $T_{c}$;
-17       $\vec{c}\leftarrow$ average document representation in $D_{c}$;
-18
-19for $d_{i}\in\mathcal{D}$ do
-20       $\mathbb{C}_{i},conf_{i}\leftarrow$ refined core classes using $\cos(\vec{d},\vec{c})$ and Eq. 6;
-21
-[…]
-23 // Text Classifier Training with Path-Based Data Augmentation;
-24 $\mathcal{D}^{\text{gen}}\leftarrow$ generate $q$ documents for each path using $\mathcal{G}$;
-25 $F\leftarrow$ train classifier with $\mathcal{D}^{\text{core}}$, $\mathcal{D}^{\text{gen}}$, and Eq. 7;
-26 Return $F$;
-Algorithm 1 TELEClass
-Table 1.
-Experiment results on Amazon-531 and DBPedia-298 datasets, evaluated by Example-F1, P@k, and MRR.
+Algorithm 1 in Appendix summarizes TELEClass.
+Table 1: Experiment results on Amazon-531 and DBPedia-298 datasets, evaluated by Example-F1, P@k, and MRR.
 The best score among zero-shot and weakly-supervised methods is boldfaced.
-“${}^{\dagger}$” indicates the numbers for these baselines are directly from previous paper (Shen et al., 2021).
+“” indicates the numbers for these baselines are directly from previous paper Shen et al.
+(2021).
 “—” means the method cannot generate a ranking of predictions and thus MRR cannot be calculated.
@@
 Example-F1 P@1 P@3 MRR Example-F1 P@1 P@3 MRR
-Zero-Shot ${}^{\dagger}$ 0.4742 0.7144 0.4610 0.6765 0.7871 0.6765
-GPT-3.5-turbo 0.5164 0.6807 0.4752 0.4816 0.5328 0.4547
-Weakly-Supervised ${}^{\dagger}$ 0.3157 0.5805 0.3115 0.1443 0.2635 0.1443
-${}^{\dagger}$ 0.2458 0.5773 0.2517 0.3047 0.5359 0.3048
-${}^{\dagger}$ 0.5431 0.7918 0.5414 0.5911 0.7712 0.8621 0.7712 0.8221
-${}^{\dagger}$ 0.5934 0.8120 0.5894 0.6332 0.8156 0.8942 0.8156 0.8762
-TELEClass 0.6330 0.8439 0.6269 0.6664 0.8684 0.9293 0.8684 0.8880
+Zero-Shot Hier-0Shot-TC 0.4742 0.7144 0.4610 0.6765 0.7871 0.6765
+ChatGPT 0.5164 0.6807 0.4752 0.4816 0.5328 0.4547
+Weakly-Supervised Hier-doc2vec 0.3157 0.5805 0.3115 0.1443 0.2635 0.1443
+WeSHClass 0.2458 0.5773 0.2517 0.3047 0.5359 0.3048
+TaxoClass-NoST 0.5431 0.7918 0.5414 0.5911 0.7712 0.8621 0.7712 0.8221
+TaxoClass 0.5934 0.8120 0.5894 0.6332 0.8156 0.8942 0.8156 0.8762
+TELEClass 0.6483 0.8505 0.6421 0.6865 0.8633 0.9351 0.8633 0.8864
 Fully-Supervised 0.8843 0.9524 0.8758 0.9085 0.9786 0.9945 0.9786 0.9826
-4. Experiments
-4.1. Experiment Setup
-4.1.1. Datasets
+4 Experiments
+4.1 Experiment Setup
+Datasets.
 We use two public datasets in different domains for evaluation.
-(1) Amazon-531 (McAuley and Leskovec, 2013) consists of Amazon product reviews and a three-layer label taxonomy of 531 product types, with 29,487 unlabeled training samples and 19,685 testing samples.
-(2) DBpedia-298 (Lehmann et al., 2015) consists of Wikipedia articles with a three-layer label taxonomy of 298 categories, with 196,665 unlabeled training samples and 49,167 testing samples.
-4.1.2. Compared Methods
-We compare the following methods on the weakly-supervised hierarchical text classification task.
-  • Hier-0Shot-TC (Yin et al., 2019) is a zero-shot approach, which utilizes a pretrained textual entailment model to iterative find the most similar class at each level for a document.
-  • GPT-3.5-turbo is a zero-shot approach that queries GPT-3.5-turbo by directly providing all classes in the prompt.
-  • Hier-doc2vec (Le and Mikolov, 2014) is a weakly-supervised approach, which first trains document and class representations in the same embedding space, and then iteratively selects the most similar class at each level.
-  • WeSHClass (Meng et al., 2019) is a weakly-supervised approach using a set of keywords for each class.
-It first generates pseudo documents to pretrain text classifiers and then performs self-training.
-  • TaxoClass and its variant TaxoClass-NoST (Shen et al., 2021) is a weakly-supervised approach that only uses the class name of each class.
-It first uses a textual entailment model with a top-down search and corpus-level comparison to select core classes, which are then used as pseudo-training data.
-We include both its full model, TaxoClass, and its variation TaxoClass-NoST that does not apply self-training on the trained classifier, which is the same as ours.
-  • TELEClass is our newly proposed weakly-supervised approach that only uses the class name for each class.
-  • Fully-Supervised is a fully-supervised baseline that uses the entire labeled training data to train the text-matching network used in TELEClass.
-4.1.3. Evaluation Metrics
-Following previous studies (Shen et al., 2021), we utilize the following evaluation metrics: Example-F1, Precision at k (P@k) for $k=1,3$, and Mean Reciprocal Rank (MRR).
-See Appendix A for definitions of these metrics.
-4.1.4. Implementation Details
-We use RoBERTa-large-MNLI as the textual entailment model for core class candidate search.
-We query GPT-3.5-turbo-0613 for both core class annotation and path-based generation.
-For taxonomy enrichment, we get the term and class name embeddings using a pre-trained BERT-base-uncased.
-The weight parameter $\alpha$ is set to 0.2, and we select top $k=20$ enriched terms for each class (c.f.
-Section 3.2).
-For core class refinement, we get the document and class embeddings using Sentence Transformer all-mpnet-base-v2 (c.f.
-Section 3.3).
-We generate $q=5$ documents with path-based generation for each class.
-The document encoder in the final classifier is initialized with BERT-base-uncased.
-We train the classifier using AdamW optimizer with a learning rate 5e-5, and the batch size is 64.
-We include prompts we used for LLM annotation and generation in Appendix B.
-4.2. Experimental Results
+(1) Amazon-531 McAuley and Leskovec (2013) consists of Amazon product reviews and a three-layer label taxonomy of 531 product types, with 29,487 unlabeled training samples and 19,685 testing samples.
+(2) DBPedia-298 Lehmann et al.
+(2015) consists of Wikipedia articles with a three-layer label taxonomy of 298 categories, with 196,665 unlabeled training samples and 49,167 testing samples.
+Compared Methods.
+We compare TELEClass with the following methods: Hier-0Shot-TC Yin et al.
+(2019), zero-shot ChatGPT (i.e., GPT-3.5-turbo), Hier-doc2vec Le and Mikolov (2014), WeSHClass Meng et al.
+(2019), TaxoClass Shen et al.
+(2021) and its variant TaxoClass-NoST that does not apply self-training.
+We also include a Fully-Supervised baseline that uses the entire labeled training data to train the text-matching network.
+See more details of these methods in Appex.
+B.1.
+The implementation details of TELEClass are in Appx.
+B.3.
+Evaluation Metrics.
+Following previous studies Shen et al.
+(2021), we utilize the following evaluation metrics: Example-F1, Precision at k (P@k) for $k=1,3$, and Mean Reciprocal Rank (MRR).
+See Appendix B.2 for detailed definitions.
+4.2 Experimental Results
 Table 1 shows the evaluation results of all the compared methods.
 We make the following observations.
-(1) Overall, TELEClass achieves the best performance among the compared strong zero-shot and weakly-supervised baselines.
-(2) By comparing with other weakly-supervised methods, we find that TELEClass significantly outperforms TaxoClass-ST, which is the strongest baseline that also does not use self-training as TELEClass.
+(1) Overall, TELEClass achieves significantly better performance than other strong zero-shot and weakly-supervised baselines.
+(2) By comparing with other weakly-supervised methods, we find that TELEClass significantly outperforms TaxoClass-ST, the strongest baseline that, like TELEClass, does not use self-training.
 This further demonstrates the effectiveness of TELEClass.
-(3) Although LLM (e.g., GPT-3.5-turbo) show superior power in many tasks, naïvely prompting GPT-3.5-turbo in the hierarchical text classification task does not yield promising performance and underperforms strong weakly-supervised text classifier by a large margin.
-We will conduct a more detailed comparison of LLM prompting for hierarchical text classification in Section 4.4.
-Table 2. Performance of TELEClass and its ablations on Amazon-531 and DBPedia-298 datasets. The best score is boldfaced.
+(3) Although LLMs (e.g., ChatGPT) show superior power in many tasks, naïvely prompting it in the hierarchical text classification task yields significantly inferior performance compared to strong weakly-supervised text classifiers.
+We conduct a more detailed comparison with LLM prompting for hierarchical text classification in Section 4.4.
+Table 2: Performance of TELEClass and its ablations on Amazon-531 and DBPedia-298.
 Methods Amazon-531 DBPedia-298
@@
 Gen-Only 0.5151 0.7477 0.5096 0.5357 0.7930 0.9421 0.7930 0.8209
-TELEClass-NoEnrich 0.6324 0.8375 0.6263 0.6730 0.6868 0.7726 0.6868 0.7182
-TELEClass-NoGen 0.6084 0.8238 0.6024 0.6361 0.8364 0.9146 0.8364 0.8599
-TELEClass 0.6330 0.8439 0.6269 0.6664 0.8684 0.9293 0.8684 0.8880
-Table 3.
-Performance comparison of zero-shot LLM prompting.
-We only report Example-F1 and P@k, because it is not straightforward to get ranking of classes predicted by LLMs for MRR calculation.
-We also report estimated cost in US dollar (as of the pricing in Feb 2024) and running time in minutes for each method on the entire test set.
-“${}^{\ddagger}$” indicates that we report the performance on a 1,000-document subset of test data.
+TELEClass-NoLLMEnrich 0.5520 0.7370 0.5463 0.5900 0.8319 0.9108 0.8319 0.8563
+TELEClass-NoCorpusEnrich 0.6143 0.8358 0.6082 0.6522 0.8185 0.8916 0.8185 0.8463
+TELEClass-NoGen 0.6449 0.8348 0.6387 0.6792 0.8494 0.9187 0.8494 0.8730
+TELEClass 0.6483 0.8505 0.6421 0.6865 0.8633 0.9351 0.8633 0.8864
+Table 3: Performance comparison of TELEClass and zero-shot LLM prompting.
 Methods Amazon-531 DBPedia-298
-Example-F1 P@1 P@3 Est. Cost Est. Time Example-F1 P@1 P@3 Est. Cost Est. Time
-GPT-3.5-turbo 0.5164 0.6807 0.4752 \$60 240 mins 0.4816 0.5328 0.4547 \$80 400 mins
-GPT-3.5-turbo (level) 0.6621 0.8574 0.6444 \$20 800 mins 0.6649 0.8301 0.6488 \$60 1,000 mins
-${}^{\ddagger}$ 0.6994 0.8220 0.6890 \$800 400 mins 0.6054 0.6520 0.5920 \$2,500 1,000 mins
-TELEClass 0.6330 0.8439 0.6269 <\$1 3 mins 0.8684 0.9293 0.8684 <\$1 7 mins
-4.3. Ablation Studies
-To better understand if each component of TELEClass contributes to the entire method, we additionally conduct ablation studies by removing part of TELEClass and reporting the performance.
-Table 2 shows the results of the following ablations of TELEClass:
-  • Gen-Only only uses the augmented documents produced by path-based LLM generation to train the final classifier.
-  • TELEClass-NoEnrich does not perform taxonomy enrichment, so directly uses the initial core classes annotated by LLM plus augmented documents to train the final classifier.
-  • TELEClass-NoGen does not use the augmented documents produced by path-based LLM generation, so only uses the refined core classes by taxonomy enrichment and embedding-based matching to train the final classifier.
+F1 P@1 P@3 F1 P@1 P@3
+ChatGPT 0.516 0.681 0.475 0.482 0.533 0.455
+ChatGPT-L 0.662 0.857 0.644 0.665 0.830 0.649
+GPT-4 0.699 0.822 0.689 0.605 0.652 0.592
+TELEClass 0.648 0.851 0.642 0.863 0.935 0.863
+4.3 Ablation Studies
+We conduct ablation studies to better understand how each component of TELEClass contributes to final performance. Table 2 shows the results of the following ablations:
+  • Gen-Only only uses the augmented documents by path-based LLM generation to train the final classifier.
+  • TELEClass-NoLLMEnrich excludes the LLM-based taxonomy enrichment component.
+  • TELEClass-NoCorpusEnrich excludes the corpus-based taxonomy enrichment component.
+  • TELEClass-NoGen excludes the augmented documents by path-based LLM generation.
 We find that the full model TELEClass achieves the overall best performance among the compared methods, showing the effectiveness of each of its components.
-Interestingly, we find that taxonomy enrichment and path-based data generation make different levels of contribution on two data sets: path-based data generation brings more improvement on Amazon-531, while taxonomy enrichment contributes more on DBpedia-298.
-The reasons could be as follows.
-Amazon-531 has a large label space and more skewed class distribution.
-We find that around 11.6% of classes are not selected as core classes in Amazon, while this number for DBpedia is 5.4%.
-This necessitates data augmentation to provide training data that covers each class in the taxonomy.
-On the other hand, although DBpedia has a smaller label space, its class names are more subtle to distinguish, which can also be demonstrated by the lower performance of zero-shot LLM prompting on DBpedia compared to Amazon-531 (c.f.
-GPT-3.5-turbo in Table 1).
-Therefore, by mining class-indicative terms, TELEClass can enrich each class with more clues to facilitate better classification.
-4.4. Comparison with Zero-Shot LLM Prompting
+First, both the LLM-based and corpus-based enrichment modules bring improvement to the performance.
+Interestingly, we find that they make different levels of contribution on the two datasets: LLM-based enrichment brings more improvement on Amazon-531 while corpus-based enrichment contributes more on DBPedia-298.
+We suspect the reasons are as follows.
+The classes in Amazon-531 are commonly seen product types that LLM can understand and enrich in a reliable manner.
+However, DBPedia-298 contains classes that are more subtle to distinguish, which can also be shown by the lower performance of zero-shot LLM prompting on DBPedia compared to Amazon-531 (c.f.
+ChatGPT in Table 1).
+Therefore, corpus-based enrichment can consolidate the meaning of each class based on corpus-specific knowledge to facilitate better classification.
+We also find that path-based LLM generation consistently improves the model performance while requiring only a few hundred queries to LLMs.
+Even Gen-Only achieves comparable performance to the strong baseline TaxoClass-NoST, demonstrating the effectiveness of this augmentation step.
+4.4 Comparison with Zero-Shot LLM Prompting
 In this section, we further compare TELEClass with zero-shot LLM prompting.
 Because it is not straightforward to get ranked predictions by LLMs, we only report Example-F1 and P@k as performance evaluations.
-Additionally, we report the estimated cost and time for each method on the entire test set.
-The cost is reported in US dollar based on the pricing in February 2024.
-The inference time is reported in minutes, and please be aware that this is just a rough estimation as the actual running time is also dependent on the server condition.
-We include the following settings to compare with TELEClass:
-  • GPT-3.5-turbo: We include all the classes in the prompt and ask GPT-3.5-turbo model to provide 3 most appropriate classes for a given document.
-  • GPT-3.5-turbo (level): We perform level-by-level prompting using GPT-3.5-turbo.
-Starting from the root node, we ask the model to return one most appropriate class for a given document, and we iteratively prompt the model with the child nodes of the selected node at each level.
-Notice that this method can only generate a path in the taxonomy, but in the actual multi-label hierarchical classification setting, the true labels may not sit in the same path in the label taxonomy.
-  • GPT-4: Similar to the first one, we include all the classes in the prompt and ask GPT-4 to provide 3 most appropriate classes for a given document.
-Given the limited budget, we only test it on a sampled subset of the test split, which contains 1,000 documents for each dataset.
+We report the zero-shot performance of ChatGPT and GPT-4 by including all the classes in the prompt and asking the LLM to provide the three most appropriate classes for a given document.
+Experiments on GPT-4 are conducted on 1000 randomly sampled test documents given the prohibitively high cost.
+We also include a level-by-level version, ChatGPT-L, which selects one most appropriate class at each level and continues to its children.
 Table 3 shows the experiment results.
-We find that TELEClass consistent outperforms all compared methods on DBpedia, while on Amazon, TELEClass underperforms GPT-3.5-turbo (level) and GPT-4 but still being comparable.
-As for the cost, once trained, TELEClass does not require an additional cost on inference and also has substantially shorter inference time.
-Prompting LLMs takes longer time and can be prohibitively expensive (e.g., using GPT-4), and the cost will scale up with an increasing size of test data.
-Also, we find that the level-by-level version of GPT-3.5-turbo consistently outperforms the naïve version, demonstrating the necessity of taxonomy structure.
-It saves the cost because of the much shorter prompts, but takes longer time due to more queries made per document.
-Table 4.
-Intermediate results on two documents, including selected core classes by different methods, true labels, and the corresponding taxonomy enrichment results.
+We find that TELEClass consistently outperforms zero shot performance of LLMs by large margins on DBPedia-298.
+On Amazon-531, while it continues to show substantial improvements over ChatGPT, it slightly underperforms ChatGPT-L and GPT-4.
+As for the cost, once trained, TELEClass does not require an additional cost on inference and also has a substantially shorter inference time.
+Prompting LLMs takes a longer time and can be prohibitively expensive, and the cost will scale up with an increasing size of test data.
+The estimated API cost for ChatGPT on the entire DBPedia test set is \$80 and for GPT-4 it is around \$2,500.
+Also, we find that the level-by-level version ChatGPT-L consistently outperforms the naïve version, demonstrating the necessity of taxonomy structure.
+Table 4: Intermediate results on two documents, including selected core classes by different methods, true labels, and the corresponding taxonomy enrichment results.
 The optimal core class in the true labels is marked with ©.
 Dataset Document Core Classes by… True Labels Corr. Enrichment
-DBPedia The Lindenhurst Memorial Library (LML) is located in Lindenhurst, New York, and is one of the fifty six libraries that are part of the Suffolk Cooperative Library System … TaxoClass: village TELEClass initial: building TELEClass refined: library library©, agent, educational institution Class: library Top Enrichment: national library, central library, collection, volumes…
+DBPedia The Lindenhurst Memorial Library (LML) is located in Lindenhurst, New York, and is one of the fifty six libraries that are part of the Suffolk Cooperative Library System … TaxoClass: village TELEClass initial: building TELEClass refined: library library©, agent, educational institution Class: library Enrichment: national library, central library, collection, volumes…
 Amazon Since mom (89 yrs young) isn’t steady on her feet, we have place these grab bars around the room.
 It gives her the stability and security she needs.
-TaxoClass: personal care, health personal care, safety TELEClass initial: daily living aids, medical supplies equipment, safety, TELEClass refined: bathroom aids safety health personal care, medical supplies equipment, bathroom aids safety© Class: bathroom aids safety Top Enrichment: seat, toilet, shower, safety, handles…
-4.5. Case studies
-To better understand the TELEClass framework, we show some intermediate results of two documents in Table 4, including the core classes selected by (1) the original TaxoClass (Shen et al., 2021) method, (2) TELEClass’s initial core classes selected by LLM, and (3) refined core classes of TELEClass.
+TaxoClass: personal care, health personal care, safety TELEClass initial: daily living aids, medical supplies equipment, safety, TELEClass refined: bathroom aids safety health personal care, medical supplies equipment, bathroom aids safety© Class: bathroom aids safety Enrichment: seat, toilet, shower, safety, handles…
+4.5 Case studies
+To better understand the TELEClass framework, we show some intermediate results of two documents in Table 4, including the core classes selected by (1) the original TaxoClass Shen et al.
+(2021) method, (2) TELEClass’s initial core classes selected by LLM, and (3) refined core classes of TELEClass.
 Besides, we also include the true labels of the documents and the taxonomy enrichment results of the corresponding core class in the table.
@@
 In the other example, TELEClass also pinpoints the most accurate core class “bathroom aids safety” while other methods can only find more general or partially relevant classes.
-5. Related Work
-5.1. Weakly-Supervised Text Classification
+5 Related Work
+5.1 Weakly-Supervised Text Classification
 Weakly-supervised text classification trains a classifier with limited guidance, aiming to reduce human efforts while maintaining high proficiency.
-Various sources of weak supervision have been explored, including distant supervision (Gabrilovich and Markovitch, 2007; Chang et al., 2008; Song and Roth, 2014) like knowledge bases, keywords (Agichtein and Gravano, 2000; Tao et al., 2018; Meng et al., 2018, 2020; Wang et al., 2021a; Zhang et al., 2021b), and heuristic rules (Ratner et al., 2016; Badene et al., 2019; Shu et al., 2020).
-Later, extremely weakly-supervised methods are proposed to solely rely on class label names for generating pseudo-labels and training classifiers.
-LOTClass (Meng et al., 2020) utilizes MLM-based PLM as a knowledge base for extracting class-indicative keywords.
-XClass (Wang et al., 2021a) extracts keywords for creating static class representations through clustering.
-PIEClass (Zhang et al., 2023a) employs PLMs’ zero-shot prompting to obtain pseudo labels with noise-robust iterative ensemble training.
-MEGClass (Kargupta et al., 2023) acquires contextualized sentence representations to capture topical information at the document level.
-WOTClass (Wang et al., 2023) focuses on mining and ranking class-indicative keywords from generated classes and extracting classes with overlapping keywords.
-5.2. Hierarchical Text Classification
+Various sources of weak supervision have been explored, including distant supervision Gabrilovich and Markovitch (2007); Chang et al.
+(2008); Song and Roth (2014) like knowledge bases, keywords Agichtein and Gravano (2000); Tao et al.
+(2018); Meng et al.
+(2018, 2020); Wang et al.
+(2021a); Zhang et al.
+(2021a), and heuristic rules Ratner et al.
+(2016); Badene et al.
+(2019); Shu et al.
+(2020).
+Later, extremely weakly-supervised methods are proposed to solely rely on class names to generate pseudo labels and train classifiers.
+LOTClass Meng et al.
+(2020) utilizes MLM-based PLM as a knowledge base for extracting class-indicative keywords.
+X-Class Wang et al.
+(2021a) extracts keywords for creating static class representations through clustering.
+PIEClass Zhang et al.
+(2023c) employs PLMs’ zero-shot prompting to obtain pseudo labels with noise-robust iterative ensemble training.
+MEGClass Kargupta et al.
+(2023) acquires contextualized sentence representations to capture topical information at the document level.
+WOTClass Wang et al.
+(2023) ranks class-indicative keywords from generated classes and extracts classes with overlapping keywords.
+5.2 Hierarchical Text Classification
 A hierarchy provides a systematic top-to-down structure with inherent semantic relations that can assist in text classification.
 Typical hierarchical text classification can be categorized into two groups: local approaches and global approaches.
-Local approaches train multiple classifiers for each node or local structures (Banerjee et al., 2019; Liu et al., 2005; Wehrmann et al., 2018).
-Global approaches, learn hierarchy structure into a single classifier through recursive regularization (Gopal and Yang, 2013), a graph neural network (GNN)-based encoder (Peng et al., 2018; Huang et al., 2019; Zhou et al., 2020; Shen et al., 2021), or a joint document label embedding space (Chen et al., 2021).
+Local approaches train multiple classifiers for each node or local structures Banerjee et al.
+(2019); Liu et al.
+(2005); Wehrmann et al.
+(2018).
+Global approaches, learn hierarchy structure into a single classifier through recursive regularization Gopal and Yang (2013), a graph neural network (GNN)-based encoder Peng et al.
+(2018); Huang et al.
+(2019); Zhou et al.
+(2020); Shen et al.
+(2021), or a joint document label embedding space Chen et al.
+(2021).
 Weak supervision is also studied for hierarchical text classification.
-WeSHClass (Meng et al., 2019) uses a few keywords or example documents per class and pretrains classifiers with pseudo documents followed by self-training.
-TaxoClass (Shen et al., 2021) follows the same setting as ours which uses the sole class name of each class as the only supervision.
+WeSHClass Meng et al.
+(2019) uses a few keywords or example documents per class and pretrains classifiers with pseudo documents followed by self-training.
+TaxoClass Shen et al.
+(2021) follows the same setting as ours which uses the sole class name of each class as the only supervision.
 It identifies core classes for each document using a textual entailment model, which is then used to train a multi-label classifier.
-Additionally, MATCH (Zhang et al., 2021c) and HiMeCat (Zhang et al., 2021a) study how to integrate associated metadata into the label hierarchy for document categorization with weak supervision.
-5.3. LLMs as Generators and Annotators
-Large language models (LLMs) have demonstrated impressive performance in many downstream tasks and are explored to help low-resource settings by synthesizing data as generators or annotators (Ding et al., 2023).
-For data generation, few-shot examples (Wang et al., 2021b) or class-conditioned prompts (Meng et al., 2022) are explored for LLM generation and the generated data can be used as pseudo-training data to further fine-tune a small model as the final classifier (Ye et al., 2022).
+Additionally, MATCH Zhang et al.
+(2021c) and HiMeCat Zhang et al.
+(2021b) study how to integrate associated metadata into the label hierarchy for document categorization with weak supervision.
+5.3 LLMs as Generators and Annotators
+Large language models (LLMs) have demonstrated impressive performance in many downstream tasks and are explored to help low-resource settings by synthesizing data as generators or annotators Ding et al.
+(2023).
+For data generation, few-shot examples Wang et al.
+(2021b) or class-conditioned prompts Meng et al.
+(2022) are explored for LLM generation and the generated data can be used as pseudo training data to further fine-tune a small model as the final classifier Ye et al.
+(2022).
 Recently, Yu et al.
-(2023) propose an attribute-aware topical text classification method that incorporates ChatGPT to generate topic-dependent attributes and topic-independent attributes to reduce topic ambiguity and increase topic diversity for generation.
-For data annotation, previous works utilize LLMs for unsupervised annotation (Feng et al., 2021), Chain-of-Thought annotation with explanation generation (He et al., 2023), and active annotation (Zhang et al., 2023b).
-6. Conclusion and Future Work
+(2023) proposed an attribute-aware topical text classification method that incorporates ChatGPT to generate topic-dependent attributes and topic-independent attributes to reduce topic ambiguity and increase topic diversity for generation.
+For data annotation, previous works utilize LLMs for unsupervised annotation Feng et al.
+(2021), Chain-of-Thought annotation with explanation generation He et al.
+(2024), and active annotation Zhang et al.
+(2023a).
+6 Conclusion and Future Work
 In this paper, we propose a new method, TELEClass, for the minimally-supervised hierarchical text classification task with two major contributions.
-First, we enrich the input label taxonomy with more class-indicative topical terms for each class, which can serve as additional features to understand the classes and facilitate classification.
+First, we enrich the input label taxonomy with LLM-generated and corpus-based class-indicative terms for each class, which can serve as additional features to understand the classes and facilitate classification.
 Second, we explore the utilization of LLMs in the hierarchical text classification in two directions: data annotation and data creation.
 On two public datasets, TELEClass can outperform existing baseline methods substantially, and we further demonstrate its effectiveness through ablation studies.
-We also conduct a comparative analysis on performance and cost for zero-shot LLM prompting for the hierarchical text classification task.
-For future works, we plan to generalize TELEClass’s idea into other low-resource text mining tasks with hierarchical label spaces, such as fine-grained entity typing.
-We also plan to explore how to extend TELEClass into a more complicated label space, such as an extremely large space with millions of labels.
-Appendix A Evaluation Metrics
+We also conduct a comparative analysis of performance and cost for zero-shot LLM prompting for the hierarchical text classification task.
+For future works, first, we plan to generalize TELEClass’s idea into other low-resource text mining tasks with hierarchical label spaces, such as fine-grained entity typing.
+We also plan to explore how to extend TELEClass into more complicated label spaces, such as an extremely large space with millions of labels.
+Limitations
+In this paper, we propose TELEClass, a novel hierarchical text classification method that uses the label taxonomy as the only supervision.
+Despite its effectiveness shown in the experiments, there is still room to improve.
+First, TELEClass utilizes the general knowledge of pretrained LLM and shows strong performance on two general domain datasets.
+However, it remains a challenge to train effective text classifiers in private domains, for which LLMs lack relevant knowledge.
+Second, in this paper, we mainly focus on acquiring high-quality pseudo labeled data while only utilizing the simplest classifier model and objective.
+It is worth studying how the proposed method can be further improved with more advanced network structure and noise-robust training objectives.
+Lastly, TELEClass assumes the availability of a large unlabeled corpus for pseudo labeling, whose performance is not fully studied if limited data is available.
+Ethics Statement
+In this work, we develop a new hierarchical text classification method and experiment on two publicly available datasets for research purposes.
+Overall, we do not foresee any significant ethical issues from our work.
+All the datasets are used as intended.
+We will open-source our code upon publication to assist future studies.
+Although showing strong language understanding capability, LLMs can sometimes generate misinformation Chen and Shu (2023) or biased contents Fang et al.
+(2024).
+In our work, instead of studying the generation ability of LLMs, we only use LLMs to generate pseudo documents for text classifier training.
+In practice, our method can be combined with any misinformation detection and debiasing methods Mattern et al.
+(2022); Chen and Shu (2021).
+Acknowledgments
+Research was supported in part by US DARPA INCAS Program No.
+HR001121C0165 and KAIROS Program No.
+FA8750-19-2-1004, National Science Foundation IIS-19-56151, and the Molecule Maker Lab Institute: An AI Research Institutes program supported by NSF under Award No.
+2019897, and the Institute for Geospatial Understanding through an Integrative Discovery Environment (I-GUIDE) by NSF under Award No.
+2118329.
+Any opinions, findings, and conclusions or recommendations expressed herein are those of the authors and do not necessarily represent the views, either expressed or implied, of DARPA or the U.S.
+Government.
+Appendix A Pseudo code of TELEClass
+Algorithm 1 summarizes the entire TELEClass framework.
+Input: A corpus $\mathcal{D}$, a label taxonomy $\mathcal{T}$, a pretrained text encoder $\mathcal{S}$, an LLM $\mathcal{G}$.
+Output: A text classifier $F$ that can classify each document into a set of classes in $\mathcal{T}$.
+1 // LLM-Enhanced Core Class Annotation;
+2 for $c\in\mathcal{C}$ do
+3       $T^{\text{LLM}}_{c}\leftarrow$ use $\mathcal{G}$ to enrich $c$ with key terms;
+4
+5for $d_{i}\in\mathcal{D}$ do
+6       $\mathbb{C}^{0}_{i}\leftarrow$ use $\mathcal{G}$ to select core classes from candidates retrieved using $\mathcal{S}$ and $T^{\text{LLM}}_{c}$;
+7
+8// Corpus-Based Taxonomy Enrichment;
+9 for $c\in\mathcal{C}$ do
+10       $D^{0}_{c}\leftarrow$ a set of roughly classified documents;
+11       for $c_{p}\in Par(c)$ do
+12             $T(c,c_{p})\leftarrow$ top terms ranked by affinity;
+13
+14      $T_{c}\leftarrow$ aggregate corpus-based and LLM-generated terms;
+15
+16// Core Class Refinement with Enriched Taxonomy;
+17 $\vec{d}\leftarrow$ document representation $\mathcal{S}(d)$;
+18 for $c\in\mathcal{C}$ do
+19       $D_{c}\leftarrow$ confident documents by matching $T_{c}$;
+20       $\vec{c}\leftarrow$ average document representation in $D_{c}$;
+21
+22for $d_{i}\in\mathcal{D}$ do
+23       $\mathbb{C}_{i},conf_{i}\leftarrow$ refined core classes using $\cos(\vec{d},\vec{c})$ and Eq. 2;
+24
+[…]
+26 // Text Classifier Training with Path-Based Data Augmentation;
+27 $\mathcal{D}^{\text{gen}}\leftarrow$ generate $q$ documents for each path using $\mathcal{G}$;
+28 $F\leftarrow$ train classifier with $\mathcal{D}^{\text{core}}$ and $\mathcal{D}^{\text{gen}}$;
+29 Return $F$;
+Algorithm 1 TELEClass
+Appendix B Experiment Setup
+B.1 Compared Methods
+We compare the following methods on the weakly-supervised hierarchical text classification task.
+  • Hier-0Shot-TC Yin et al. (2019) is a zero-shot approach, which utilizes a pretrained textual entailment model to iterative find the most similar class at each level for a document.
+  • Hier-doc2vec Le and Mikolov (2014) is a weakly-supervised approach, which first trains document and class representations in a same embedding space, and then iteratively selects the most similar class at each level.
+  • WeSHClass Meng et al.
+(2019) is a weakly-supervised approach using a set of keywords for each class.
+It first generates pseudo documents to pretrain text classifiers and then performs self-training.
+  • TaxoClass Shen et al.
+(2021) is a weakly-supervised approach that only uses the class name of each class.
+It first uses a textual entailment model with a top-down search and corpus-level comparison to select core classes, which are then used as pseudo training data.
+We include both its full model, TaxoClass, and its variation TaxoClass-NoST that does not apply self-training on the trained classifier, which is the same as TELEClass.
+B.2 Evaluation Metrics
 Let $\mathbb{C}^{\text{true}}_{i}$ and $\mathbb{C}^{\text{pred}}_{i}$ denote the set of true labels and the set of predicted labels for document $d_{i}\in\mathcal{D}$, respectively.
@@
 The evaluation metrics are defined as follows:
-  • Example-F1 (Sørensen, 1948), which is also called micro-Dice coefficient, evaluates the multi-label classification results without ranking,
+  • Example-F1 Sørensen (1948), which is also called micro-Dice coefficient, evaluates the multi-label classification results without ranking,
 […]
@@
   • Mean Reciprocal Rank, or MRR, is another ranking-based metric, which evaluates the multi-label predictions based on the inverse of true labels’ ranks within predicted classes,
-    (10) $\text{MRR}=\frac{1}{|\mathcal{D}|}\sum_{d_{i}\in\mathcal{D}}\frac{1}{|\mathbb{C}_{i}^{\text{true}}|}\sum_{c_{j}\in\mathbb{C}_{i}^{\text{true}}}\frac{1}{\min\{k|c_{j}\in\mathbb{C}^{\text{pred}}_{i,k}\}}.$
-Appendix B Prompts for LLM
+    $\text{MRR}=\frac{1}{|\mathcal{D}|}\sum_{d_{i}\in\mathcal{D}}\frac{1}{|\mathbb{C}_{i}^{\text{true}}|}\sum_{c_{j}\in\mathbb{C}_{i}^{\text{true}}}\frac{1}{\min\{k|c_{j}\in\mathbb{C}^{\text{pred}}_{i,k}\}}.$ (5)
+B.3 Implementation Details
+We use Sentence Transformer Reimers and Gurevych (2019) all-mpnet-base-v2 as the text encoder for the similarity measure in Section 3.1 and Section 3.3.
+We query GPT-3.5-turbo-0125 for LLM-based taxonomy enrichment, core class annotation, and path-based generation.
+For corpus-based taxonomy enrichment, we get the term and class name embeddings using a pre-trained BERT-base-uncased Devlin et al.
+(2019), we select top $k=20$ enriched terms for each class (c.f.
+Section 3.2).
+We generate $q=5$ documents with path-based generation for each class.
+The document encoder in the final classifier is initialized with BERT-base-uncased for a fair comparison with the baselines.
+We train the classifier using AdamW optimizer with a learning rate 5e-5, and the batch size is 64.
+The experiments are run on one NVIDIA RTX A6000 GPU.
+Appendix C Prompts for LLM
+  • LLM-based enrichment for Amazon-531
+    Instruction: [Target Class] is a product class in Amazon and is the subclass of [Parent Class].
+Please generate 10 additional key terms about the [Target Class] that is relevant to [Target Class] but irrelevant to [Sibling Classes].
+Please split the additional key terms using commas.
+  • LLM-based enrichment for DBPedia-298
+    Instruction: [Target Class] is a article category of Wikipedia articles and is the subclass of [Parent Class].
+Please generate 10 additional key terms about the [Target Class] that is relevant to [Target Class] but irrelevant to [Sibling Classes].
+Please split the additional key terms using commas.
   • Core class annotation for Amazon-531
-    Instruction: You will be provided with an Amazon product review, and your task is to select its product types from the following categories: [Candidate Classes].
+    Instruction: You will be provided with an Amazon product review, and please select its product types from the following categories: [Candidate Classes].
 Just give the category names as shown in the provided list.
@@
   • Core class annotation for DBPedia-298
-    Instruction: You will be provided with a Wikipedia article describing an entity at the beginning, and your task is to select its types from the following categories: [Candidate Classes].
+    Instruction: You will be provided with a Wikipedia article describing an entity at the beginning, and please select its types from the following categories: [Candidate Classes].
 Just give the category names as shown in the provided list.

### v3

@@
 TELEClass: Taxonomy Enrichment and LLM-Enhanced Hierarchical Text Classification with Minimal Supervision
-Abstract
-Hierarchical text classification aims to categorize each document into a set of classes in a label taxonomy.
+Abstract.
+Hierarchical text classification aims to categorize each document into a set of classes in a label taxonomy, which is a fundamental web text mining task with broad applications such as web content analysis and semantic indexing.
 Most earlier works focus on fully or semi-supervised methods that require a large amount of human annotated data which is costly and time-consuming to acquire.
-To alleviate human efforts, in this paper, we work on hierarchical text classification with the minimal amount of supervision: using the sole class name of each node as the only supervision.
-Recently, large language models (LLM) show competitive performance on various tasks through zero-shot prompting, but this method performs poorly in the hierarchical setting, because it is ineffective to include the large and structured label space in a prompt.
+To alleviate human efforts, in this paper, we work on hierarchical text classification with a minimal amount of supervision: using the sole class name of each node as the only supervision.
+Recently, large language models (LLM) have shown competitive performance on various tasks through zero-shot prompting, but this method performs poorly in the hierarchical setting because it is ineffective to include the large and structured label space in a prompt.
 On the other hand, previous weakly-supervised hierarchical text classification methods only utilize the raw taxonomy skeleton and ignore the rich information hidden in the text corpus that can serve as additional class-indicative features.
-To tackle the above challenges, we propose TELEClass, Taxonomy Enrichment and LLM-Enhanced weakly-supervised hierarchical text Classification, which (1) automatically enriches the label taxonomy with class-indicative terms to facilitate classifier training and (2) utilizes LLMs for both data annotation and creation tailored for the hierarchical label space.
-Experiments show that TELEClass can outperform previous weakly-supervised methods and LLM-based zero-shot prompting methods on two public datasets.
-TELEClass: Taxonomy Enrichment and LLM-Enhanced Hierarchical Text Classification with Minimal Supervision
-Yunyi Zhang1, Ruozhen Yangthanks: Equal contribution1, Xueqiang Xu11footnotemark: 11, Rui Li11footnotemark: 12, Jinfeng Xiao1, Jiaming Shen3, Jiawei Han1, 1University of Illinois Urbana-Champaign, 2University of Science and Technology of China, 3Google DeepMind, yzhan238@illinois.edu
-1 Introduction
-Hierarchical text classification, aiming to classify documents into one or multiple classes in a label taxonomy, is an essential and fundamental task in text mining and natural language processing.
+To tackle the above challenges, we propose TELEClass, Taxonomy Enrichment and LLM-Enhanced weakly-supervised hierarchical text Classification, which combines the general knowledge of LLMs and task-specific features mined from an unlabeled corpus.
+TELEClass automatically enriches the raw taxonomy with class-indicative features for better label space understanding and utilizes novel LLM-based data annotation and generation methods specifically tailored for the hierarchical setting.
+Experiments show that TELEClass can significantly outperform previous baselines while achieving comparable performance to zero-shot prompting of LLMs with drastically less inference cost.
+Weakly-Supervised Text Classification, Hierarchical Text Classification, Taxonomy Enrichment, Large Language Model
+Equal Contribution.
+journalyear: 2025copyright: ccconference: Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Proceedings of the ACM Web Conference 2025 (WWW ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3696410.3714940isbn: 979-8-4007-1274-6/25/04ccs: Information systems Data miningccs: Computing methodologies Natural language processing
+1. Introduction
+Hierarchical text classification, aiming to classify documents into one or multiple classes in a label taxonomy, is a fundamental task in web text mining and NLP.
 Compared with standard text classification where label space is flat and relatively small (e.g., less than 20 classes), hierarchical text classification is more challenging given the larger and more structured label space and the existence of fine-grained and long-tail classes.
-Most earlier works tackle this task in fully supervised You et al.
-(2019); Huang et al.
-(2019); Chen et al.
-(2021) or semi-supervised settings Berthelot et al.
-(2019); Gururangan et al.
-(2019), and different models are proposed to learn from a substantial amount of human-labeled data.
+Hierarchical text classification has broad applications such as web content organization (Dumais and Chen, 2000), semantic indexing (Li et al., 2019; Kang et al., 2024), and query classification (Cao et al., 2009; He et al., 2023, 2024b).
+Recent studies also show that hierarchically structured text (Sarthi et al., 2024; Edge et al., 2024) and document-level tagging (Poliakov and Shvai, 2024) can improve retrieval-augmented generation for large language models.
+Figure 1. An example document tagged with 3 classes. We automatically enrich each node with class-indicative terms and utilize LLMs to facilitate classification.
+The key challenge of hierarchical text classification is how to understand the large structured label space to distinguish the semantics of similar classes.
+Most earlier works tackle this task in fully supervised (You et al., 2019; Huang et al., 2019; Chen et al., 2021) or semi-supervised settings (Berthelot et al., 2019; Gururangan et al., 2019), and different models are proposed to learn from a substantial amount of human-labeled data.
 However, acquiring human annotation is often costly, time-consuming, and not scalable.
-Figure 1: An example document tagged with 3 classes. We automatically enrich each node with class-indicative terms and utilize LLMs to facilitate classification.
-Recently, large language models (LLM) such as GPT-4 OpenAI (2023) and Claude 3 Anthropic (2024) have demonstrated strong performance in flat text classification through zero-shot or few-shot prompting Sun et al.
-(2023).
-However, applying LLMs in hierarchical settings, with large and structured label spaces, remains challenging.
-Directly including hundreds of classes in prompts is ineffective, leading to structural information loss, increased computational costs, and diminished clarity for LLMs at distinguishing class-specific information.
+Recently, large language models (LLM) such as GPT-4 (OpenAI, 2023) and Claude 3 (Anthropic, 2024) have demonstrated strong performance in flat text classification (Sun et al., 2023).
+However, applying LLMs in hierarchical settings remains challenging (Yu et al., 2023a).
+Directly including hundreds of classes in prompts is ineffective and inefficient, leading to structural information loss, diminished clarity for LLMs at distinguishing class-specific information, and prohibitively expensive inference cost given the long prompt for each test document.
 Along another line of research, Meng et al.
-(2019) propose to train a moderate-size text classifier by utilizing a small set of keywords or labeled documents for each class.
+(2019) propose to train a moderate-size text classifier by utilizing a small set of keywords or labeled documents for each class and a large unlabeled corpus.
 However, compiling keyword lists for hundreds of classes and obtaining representative documents for each specific and niche category still demand significant human efforts.
@@
 Specifically, they introduce TaxoClass which generates pseudo labels with a textual entailment model for classifier training.
-However, this method overlooks additional class-relevant features in the text and suffers from the unreliable pseudo label selection due to domain shift of the entailment model.
-In this study, we advance minimally supervised hierarchical text classification by enriching the weak supervision with additional textual clues and leveraging LLMs tailored for the hierarchical label structure.
-First, we enrich the existing label taxonomy with class-specific terms derived from two sources: LLM generation and automated extraction from the corpus.
+However, this method overlooks additional class-relevant features in the corpus that could be helpful for label space understanding.
+It also suffers from the unreliable pseudo label selection because the entailment model is not trained to compare which class is more relevant to the document.
+In this study, we advance minimally supervised hierarchical text classification by taking the advantage of both LLMs’ text understanding ability and task-specific knowledge of the unlabeled text corpus.
+First, we tackle the challenge of label space understanding by enriching the label taxonomy with class-specific terms derived from two sources: LLM generation and automated extraction from the corpus.
 For example, the “conditioner” class in Figure 1 is enriched with key terms like “moisture” and “soft hair”, which distinguish it from other classes.
 These terms enhance the supervision signal by combining knowledge from LLMs and text corpus and improve the pseudo label quality for classifier training.
-Second, we further explore the application of LLMs in hierarchical text classification from two perspectives.
-Specifically, we enhance LLM annotation efficiency and effectiveness through a taxonomy-guided candidate search and optimize LLM document generation to create more precise pseudo data by using taxonomy paths.
+Second, we improve LLMs’ ability in hierarchical text classification from two perspectives: we enhance LLM annotation efficiency and effectiveness through a taxonomy-guided candidate search and also optimize LLM-based document generation to create more precise pseudo data by using taxonomy paths.
 Leveraging the above ideas, we introduce TELEClass: Taxonomy Enrichment and LLM-Enhanced weakly-supervised hierarchical text Classification.
@@
 (2) Corpus-Based Taxonomy Enrichment, where we analyze the taxonomy structure to additionally identify class-indicative topical terms through semantic and statistical analysis on the corpus.
-(3) Core Class Refinement with Enriched Taxonomy, where we embed documents and classes based on the enriched label taxonomy and refine the initially selected core classes by identifying the most similar classes for each document.
+(3) Core Class Refinement with Enriched Taxonomy, where we embed documents and classes based on the enriched label taxonomy and refined the initially selected core classes by identifying the most similar classes for each document.
 (4) Text Classifier Training with Path-Based Data Augmentation, where we sample label paths from the taxonomy and guide the LLM to generate pseudo documents most accurately describing these fine-grained classes.
 Finally, we train the text classifier on two types of pseudo labels, the core classes and the generated data, with a simple text matching network and multi-label training strategy.
-The contributions of this paper are summarized as follows: (1) We propose TELEClass, a new method for minimally supervised hierarchical text classification, which requires only the class names of the label taxonomy as supervision to train a multi-label text classifier.
-(2) We propose to enrich the label taxonomy with class-indicative terms, based on which we utilize an embedding-based document-class matching method to improve the pseudo label quality.
-(3) We study two ways of adopting large language models to help hierarchical text classifier training, which can improve the pseudo label quality and solve the data scarcity issue for fine-grained classes.
-(4) Extensive experiments on two datasets show that TELEClass can outperform strong weakly-supervised hierarchical text classification methods and zero-shot LLM prompting methods.
-2 Problem Definition
-The minimally-supervised hierarchical text classification task aims to train a text classifier that can categorize each document into multiple nodes on a label taxonomy by using the name of each node as the only supervision Shen et al.
-(2021).
+The contributions of this paper are summarized as follows:
+  • We propose TELEClass, a new method for minimally supervised hierarchical text classification, which requires only the class names of the label taxonomy as supervision to train a multi-label text classifier.
+  • We propose to enrich the label taxonomy with class-indicative terms, based on which we utilize an embedding-based document-class matching method to improve the pseudo label quality.
+  • We study two ways of adopting large language models to hierarchical text classification, which can improve the pseudo label quality and solve the data scarcity issue for fine-grained classes.
+  • Experiments on two datasets show that TELEClass can significantly outperform zero-shot and weakly-supervised hierarchical text classification baselines, while also achieving comparable performance to GPT-4 with drastically less inference cost.
+[^1] Our code and datasets can be found at: https://github.com/yzhan238/TELEClass
+2. Problem Definition
+The minimally-supervised hierarchical text classification task aims to train a text classifier that can categorize each document into multiple nodes on a label taxonomy by using the name of each node as the only supervision (Shen et al., 2021).
 For example, in Figure 1, the input document is classified as “hair care”, “shampoo”, and “scalp treatment”.
 […]
-Figure 2: Overview of the TELEClass framework.
+Figure 2. Overview of the TELEClass framework.
 We assume the label taxonomy to be a DAG instead of a tree, because it aligns better with real applications as one node can have multiple parents with different meanings.
-It is also more challenging because the classifier needs to assign a document multiple labels in different levels and paths.
-3 Methodology
+Therefore, the classifier needs to assign a document multiple labels in different levels and paths.
+Also, this setting does not restrict the prediction to be one label per level.
+In this way, the output can consist of labels from multiple paths or not necessarily contain a leaf node if the document is not specifically discussing a fine-grained topic.
+Such a setting is more flexible to fit various applications and also a more challenging research task.
+3. Methodology
 In this section, we will introduce TELEClass consisting of the following modules: (1) LLM-enhanced core class annotation, (2) corpus-based taxonomy enrichment, (3) core class refinement with enriched taxonomy, and (4) text classifier training with path-based data augmentation.
 Figure 2 shows an overview of TELEClass.
-3.1 LLM-Enhanced Core Class Annotation
-Inspired by previous studies, we first tag each document with its “core classes”, which are defined as a set of classes that can describe the document most accurately Shen et al.
-(2021).
+3.1. LLM-Enhanced Core Class Annotation
+Inspired by previous studies, we first tag each document with its “core classes”, which are defined as a set of classes that can describe the document most accurately (Shen et al., 2021).
 This process also mimics the process of human performing hierarchical text classification: first select a set of most essential classes for the document and then trace back to their relevant classes to complete the labeling.
@@
 With this enriched set of terms for each class, we define the similarity score between a document $d$ and a class $c$ as the maximum cosine similarity with the key terms:
-$sim(c,d)=\max_{t\in T^{\text{LLM}}_{c}}\cos(\vec{t},\vec{d}),\vspace{-0.5em}$ (1)
-where $\vec{\circ}$ denotes a vector representation by a pre-trained semantic encoder (e.g., Sentence Transformer Reimers and Gurevych (2019)).
+(1) $\small sim(c,d)=\max_{t\in T^{\text{LLM}}_{c}}\cos(\vec{t},\vec{d}),$
+where $\vec{\circ}$ denotes a vector representation by a pre-trained semantic encoder (e.g., Sentence Transformer (Reimers and Gurevych, 2019)).
 This newly defined similarity measure is then used for candidate core class selection.
@@
 Finally, we instruct an LLM to select the core classes for each document from the selected candidates, which produces an initial set of core classes (denoted as $\mathbb{C}^{0}_{i}$) for each document $d_{i}\in\mathcal{D}$.
-3.2 Corpus-Based Taxonomy Enrichment
+3.2. Corpus-Based Taxonomy Enrichment
 In the previous step, we enrich the raw taxonomy structure with LLM-generated key terms, which are derived from the general knowledge of LLMs but may not accurately reflect the corpus-specific knowledge.
@@
 We first collect a set of relevant documents $D^{0}_{c}\subset\mathcal{D}$ for each class $c$, which contains all the documents whose initial core classes contain $c$ or its descendants.
-Then, inspired by Tao et al.
-(2018); Zhang et al.
-(2023b), we consider the following three factors for class-indicative term selection: (1) Popularity: a class-indicative term $t$ of a class $c$ should be frequently mentioned by its relevant documents, which is quantified by the document frequency $\log(1+df(t,D^{0}_{c}))$.
-(2) Distinctiveness: a class-indicative term $t$ for a class $c$ should be infrequent in its siblings, which is quantified as the softmax of $BM25(t,D^{0}_{c})$ relevance scores over the set of siblings, $Sib(c,c_{p})$.
-(3) Semantic similarity: a class-indicative term should also be semantically similar to the class name, which is quantified as the cosine similarity between their embeddings derived from a pre-trained encoder (e.g., BERT Devlin et al.
-(2019)).
+Then, inspired by (Tao et al., 2018; Zhang et al., 2023c), we consider the following three factors for class-indicative term selection and adapt them to the hierarchical setting.
+  • Popularity: a class-indicative term $t$ of a class $c$ should be frequently mentioned by its relevant documents, which is quantified by the log normalization of its document frequency,
+    (2) $pop(t,c)=\log(1+df(t,D^{0}_{c})),$
+    where $df(t,D)$ stands for the number of documents in $D$ that mention $t$.
+  • Distinctiveness: a class-indicative term $t$ for a class $c$ should be infrequent in its siblings, which is quantified as the softmax of BM25 relevance function (Robertson and Zaragoza, 2009) over the set of siblings,
+    (3) $dist(t,c,c_{p})=\frac{\exp(BM25(t,D^{0}_{c}))}{1+\sum_{c^{\prime}\in Sib(c,c_{p})}\exp(BM25(t,D^{0}_{c^{\prime}}))}.$
+  • Semantic similarity: a class-indicative term $t$ should also be semantically similar to the class name of $c$, which is quantified as the cosine similarity between their embeddings derived from a pre-trained encoder (e.g., BERT (Devlin et al., 2019)), denoted as $sem(c,t)$.
 Finally, we define the affinity score between a term $t$ and a class $c$ corresponding to parent $p$ to be the geometric mean of the above scores, denoted as $\text{aff}(t,c,c_{p})$.
-To enrich the taxonomy, we first apply a phrase mining tool, AutoPhrase Shang et al.
-(2018), to mine quality single-token and multi-token phrases from the corpus as candidate terms [^2] Our method is flexible with any kinds of phrase mining methods like Gu et al.
+To enrich the taxonomy, we first apply a phrase mining tool, AutoPhrase (Shang et al., 2018), to mine quality single-token and multi-token phrases from the corpus as candidate terms [^3] Our method is flexible with any kinds of phrase mining methods like Gu et al.
 (2021)..
 Then, for each class $c$ and each of its parents $c_{p}$, we select the top-$k$ terms with the highest affinity scores with $c$ corresponding to $p$, denoted as $T(c,c_{p})$.
-Then, we take the union of these corpus-based terms together with the LLM-generated terms in the previous step to get the final enriched class-indicative terms for class $c$, $T_{c}=\left(\bigcup_{c_{p}\in Par(c)}T(c,c_{p})\right)\bigcup T^{\text{LLM}}_{c}$.
-3.3 Core Class Refinement with Enriched Taxonomy
+Then, we take the union of these corpus-based terms together with the LLM-generated terms in the previous step to get the final enriched class-indicative terms for class $c$,
+(4) $T_{c}=\left(\bigcup_{c_{p}\in Par(c)}T(c,c_{p})\right)\bigcup T^{\text{LLM}}_{c}.$
+3.3. Core Class Refinement with Enriched Taxonomy
 With the enriched class-indicative terms for each class, we propose to further utilize them to refine the initial core classes.
 In this paper, we adopt an embedding-based document-class matching method.
-Unlike previous methods in flat text classification Wang et al.
-(2021a) that use keyword-level embeddings to estimate document and class representations, here, we are able to define class representations directly based on document-level embeddings thanks to the rough class assignments we created in the core class annotation step (c.f.
+Unlike previous methods in flat text classification (Wang et al., 2021a) that use keyword-level embeddings to estimate document and class representations, here, we are able to define class representations directly based on document-level embeddings thanks to the rough class assignments we created in the core class annotation step (c.f.
 Sec.
 3.1).
-To obtain document representations, we utilize a pre-trained Sentence Transformer model Reimers and Gurevych (2019) to encode the entire document, which we denote as $\vec{d}$.
+To obtain document representations, we utilize a pre-trained Sentence Transformer model (Reimers and Gurevych, 2019) to encode the entire document, which we denote as $\vec{d}$.
 Then, for each class $c$, we identify a subset of its assigned documents that explicitly mention at least one of the class-indicative keywords and thus most confidently belong to this class, $D_{c}=\{d\in D^{0}_{c}|\exists w\in T_{c},w\in d\}$.
@@
 After that, we treat the classes ranked above this position as this document’s refined core classes $\mathbb{C}_{i}$, and the corresponding similarity gap as the confidence estimation $conf_{i}$.
-$\begin{split}conf_{i}=\text{diff}^{\,i}(m_{i}),\quad\mathbb{C}_{i}&=\{c^{i}_{1},\dots,c^{i}_{m_{i}}\},\\ m_{i}=\operatorname*{arg\,max}_{j\in\{1,\dots,|\mathcal{C}|-1\}}&\text{diff}^{\,i}(j).\end{split}$ (2)
-Finally, we select the top 75% of documents $d_{i}$ and their refined core classes with the highest confidence scores $conf_{i}$, denoted as $\mathcal{D}^{\text{core}}$.
-3.4 Text Classifier Training with Path-Based Data Augmentation
+(5) $\begin{split}conf_{i}=\text{diff}^{\,i}(m_{i}),\quad\mathbb{C}_{i}&=\{c^{i}_{1},\dots,c^{i}_{m_{i}}\},\\ m_{i}=\operatorname*{arg\,max}_{j\in\{1,\dots,|\mathcal{C}|-1\}}&\text{diff}^{\,i}(j).\end{split}$
+Finally, we select top 75% of documents $d_{i}$ and their refined core classes with the highest confidence scores $conf_{i}$, denoted as $\mathcal{D}^{\text{core}}$.
+3.4. Text Classifier Training with Path-Based Data Augmentation
 The final step of TELEClass is to train a hierarchical text classifier using the confident refined core classes.
@@
 By adding the generated documents to the pseudo-labeled data, we can ensure that each class of the taxonomy will be a positive class of at least $q$ documents.
+Because we generate a small constant number of documents for each label path, it also does not affect the distribution of the frequent classes.
 Moreover, we use a path instead of a single class to guide the LLM generation, because the meaning of lower-level classes is often conditioned on their parents.
@@
 To promote data diversity, we make one LLM query for each path and ask it to generate $q$ diverse documents.
+We denote the generated documents as $\mathcal{D}^{\text{gen}}$.
 Appx.
-C shows the prompts we used.
-We denote the generated documents as $\mathcal{D}^{\text{gen}}$.
+B shows the prompts we used.
 Now, with two sets of data, the pseudo-labeled documents $\mathcal{D}^{\text{core}}$ and LLM-generated documents $\mathcal{D}^{\text{gen}}$, we are ready to introduce the classifier architecture and the training process.
 Classifier architecture.
-We use a simple text matching network Shen et al.
-(2021) as our model architecture, which includes a document encoder initialized with a pre-trained BERT-base model Devlin et al.
-(2019) and a log-bilinear matching network.
+We use a simple text matching network similar to (Shen et al., 2021) as our model architecture, which includes a document encoder initialized with a pre-trained BERT-base model (Devlin et al., 2019) and a log-bilinear matching network.
 Class representations are initialized by class name embeddings (c.f.
 Sec.
-3.2).
+3.2) and are detached from the encoder model, so only the embeddings will be updated without back-propagation to the backbone model.
+Formally, the classifier predicts the probability of document $d_{i}$ belonging to class $c_{j}$:
+$\small p(c_{j}|d_{i})=\mathcal{P}(y_{j}=1|d_{i})=\sigma(\exp{(\mathbf{c}_{j}^{T}\mathbf{W}\mathbf{d}_{i})}),$
+where $\sigma$ is the sigmoid function, $\mathbf{W}$ is a learnable interaction matrix, and $\mathbf{c}_{j}$ and $\mathbf{d}_{i}$ are the encoded class and document.
 Training process.
-For each document with refined core classes, we construct its positive classes as the union of its refined core classes and their ancestors in the label taxonomy, and its negative classes are the ones that are not positive classes or descendants of any core class.
+For each document $d_{i}$ tagged with core classes $\mathbb{C}_{i}$, we construct its positive classes $\mathbb{C}^{\text{core}}_{i,+}$ as the union of its core classes and their ancestors in the label taxonomy, and its negative classes $\mathbb{C}^{\text{core}}_{i,-}$ are the ones that are not positive classes or descendants of any core class.
 This is because the ancestors of confident core classes are also likely to be true labels, and the descendants may not all be negative given that the automatically generated core classes are not optimal.
-For the LLM-generated documents, we are more confident in their labels, so we simply treat all the classes in the corresponding label path as positive classes and all other classes as negative.
-We train the text-matching model with the standard binary cross-entropy loss.
+Formally,
+$\small\mathbb{C}^{\text{core}}_{i,+}$ $=\mathbb{C}_{i}\cup\left(\cup_{c\in\mathbb{C}_{i}}Anc(c)\right),$
+$\mathbb{C}^{\text{core}}_{i,-}$ $=\mathcal{C}-\mathbb{C}^{\text{core}}_{i,+}-\cup_{c\in\mathbb{C}_{i}}Des(c),$
+where $Anc(c)$ and $Des(c)$ denote the set of ancestors and descendants and class $c$, respectively.
+For the LLM-generated documents, we are more confident in its pseudo labels, so we simply treat all the classes in the corresponding path as positive classes and all other classes as negative,
+$\small\mathbb{C}^{\text{gen}}_{p,+}=\mathbb{C}_{p},\quad\mathbb{C}^{\text{gen}}_{p,-}=\mathcal{C}-\mathbb{C}_{p}.$
+Then, we train a multi-label classifier with the binary cross entropy loss (BCE):
+$\mathcal{L}^{\text{core}}=-\sum_{d_{i}\in\mathcal{D}^{\text{core}}}\left(\sum_{c_{j}\in\mathbb{C}^{\text{core}}_{i,+}}\log{p(c_{j}|d_{i})}+\sum_{c_{j}\in\mathbb{C}^{\text{core}}_{i,-}}\log{(1-p(c_{j}|d_{i}))}\right)$
+$\mathcal{L}^{\text{gen}}=-\sum_{d^{p}_{i}\in\mathcal{D}^{\text{gen}}}\left(\sum_{c_{j}\in\mathbb{C}^{\text{gen}}_{p,+}}\log{p(c_{j}|d^{p}_{i})}+\sum_{c_{j}\in\mathbb{C}^{\text{gen}}_{p,-}}\log{(1-p(c_{j}|d^{p}_{i}))}\right)$
+(6) $\small\mathcal{L}=\mathcal{L}^{\text{core}}+\frac{|\mathcal{D}^{\text{core}}|}{|\mathcal{D}^{\text{gen}}|}\cdot\mathcal{L}^{\text{gen}}$
 The loss terms of two sets of data are weighted by their relative size, $\frac{|\mathcal{D}^{\text{core}}|}{|\mathcal{D}^{\text{gen}}|}$.
-Notice that we do not continue training the classifier with self-training that is commonly used in previous studies Meng et al.
-(2019); Shen et al.
-(2021).
+Notice that we do not continue training the classifier with self-training that is commonly used in previous studies (Meng et al., 2019; Shen et al., 2021).
 Using self-training may further improve the model performance, which we leave for future exploration.
 Algorithm 1 in Appendix summarizes TELEClass.
-Table 1: Experiment results on Amazon-531 and DBPedia-298 datasets, evaluated by Example-F1, P@k, and MRR.
+Table 1. Datasets overview.
+Dataset # unlabeled train # test # labels
+Amazon-531 29,487 19,685 531
+DBPedia-298 196,665 49,167 298
+Table 2.
+Experiment results on Amazon-531 and DBPedia-298 datasets, evaluated by Example-F1, P@k, and MRR.
 The best score among zero-shot and weakly-supervised methods is boldfaced.
-“” indicates the numbers for these baselines are directly from previous paper Shen et al.
-(2021).
+“” indicates the numbers for these baselines are directly from previous paper (Shen et al., 2021).
 “—” means the method cannot generate a ranking of predictions and thus MRR cannot be calculated.
@@
 Fully-Supervised 0.8843 0.9524 0.8758 0.9085 0.9786 0.9945 0.9786 0.9826
-4 Experiments
-4.1 Experiment Setup
-Datasets.
-We use two public datasets in different domains for evaluation.
-(1) Amazon-531 McAuley and Leskovec (2013) consists of Amazon product reviews and a three-layer label taxonomy of 531 product types, with 29,487 unlabeled training samples and 19,685 testing samples.
-(2) DBPedia-298 Lehmann et al.
-(2015) consists of Wikipedia articles with a three-layer label taxonomy of 298 categories, with 196,665 unlabeled training samples and 49,167 testing samples.
-Compared Methods.
-We compare TELEClass with the following methods: Hier-0Shot-TC Yin et al.
-(2019), zero-shot ChatGPT (i.e., GPT-3.5-turbo), Hier-doc2vec Le and Mikolov (2014), WeSHClass Meng et al.
-(2019), TaxoClass Shen et al.
-(2021) and its variant TaxoClass-NoST that does not apply self-training.
-We also include a Fully-Supervised baseline that uses the entire labeled training data to train the text-matching network.
-See more details of these methods in Appex.
-B.1.
-The implementation details of TELEClass are in Appx.
-B.3.
-Evaluation Metrics.
-Following previous studies Shen et al.
-(2021), we utilize the following evaluation metrics: Example-F1, Precision at k (P@k) for $k=1,3$, and Mean Reciprocal Rank (MRR).
-See Appendix B.2 for detailed definitions.
-4.2 Experimental Results
-Table 1 shows the evaluation results of all the compared methods.
-We make the following observations.
-(1) Overall, TELEClass achieves significantly better performance than other strong zero-shot and weakly-supervised baselines.
-(2) By comparing with other weakly-supervised methods, we find that TELEClass significantly outperforms TaxoClass-ST, the strongest baseline that, like TELEClass, does not use self-training.
-This further demonstrates the effectiveness of TELEClass.
-(3) Although LLMs (e.g., ChatGPT) show superior power in many tasks, naïvely prompting it in the hierarchical text classification task yields significantly inferior performance compared to strong weakly-supervised text classifiers.
-We conduct a more detailed comparison with LLM prompting for hierarchical text classification in Section 4.4.
-Table 2: Performance of TELEClass and its ablations on Amazon-531 and DBPedia-298.
+4. Experiments
+4.1. Experiment Setup
+4.1.1. Datasets
+We use two public datasets in different domains for evaluation. Table 1 shows the data statistics.
+  • Amazon-531 (McAuley and Leskovec, 2013) consists of Amazon product reviews and a three-layer label taxonomy of product types.
+  • DBPedia-298 (Lehmann et al., 2015) consists of Wikipedia articles with a three-layer label taxonomy of its categories.
+4.1.2. Compared Methods
+We compare the following methods on the weakly-supervised hierarchical text classification task.
+  • Hier-0Shot-TC (Yin et al., 2019) is a zero-shot approach, which utilizes a pretrained textual entailment model to iterative find the most similar class at each level for a document.
+  • GPT-3.5-turbo is a zero-shot approach that queries GPT-3.5-turbo by directly providing all classes in the prompt.
+  • Hier-doc2vec (Le and Mikolov, 2014) is a weakly-supervised approach, which first trains document and class representations in a same embedding space, and then iteratively selects the most similar class at each level.
+  • WeSHClass (Meng et al., 2019) is a weakly-supervised approach using a set of keywords for each class.
+It first generates pseudo documents to pretrain text classifiers and then performs self-training.
+  • TaxoClass (Shen et al., 2021) is a weakly-supervised approach that only uses the class name of each class.
+It first uses a textual entailment model with a top-down search and corpus-level comparison to select core classes, which are then used as pseudo training data.
+We include both its full model, TaxoClass, and its variation TaxoClass-NoST that does not apply self-training on the trained classifier, which is the same as TELEClass.
+  • TELEClass is our newly proposed weakly-supervised approach that only uses the class name for each class.
+  • Fully-Supervised is a fully-supervised baseline that uses the entire labeled training data to train the text matching network used in TELEClass.
+4.1.3. Evaluation Metrics
+Following previous studies (Shen et al., 2021), we utilize the following evaluation metrics: Example-F1, Precision at k (P@k) for $k=1,3$, and Mean Reciprocal Rank (MRR).
+See Appendix C for detailed definitions.
+4.1.4. Implementation Details
+We use Sentence Transformer (Reimers and Gurevych, 2019) all-mpnet-base-v2 as the text encoder for the similarity measure in Section 3.1 and Section 3.3.
+We query GPT-3.5-turbo-0125 for LLM-based taxonomy enrichment, core class annotation, and path-based generation.
+For corpus-based taxonomy enrichment, we get the term and class name embeddings using a pre-trained BERT-base-uncased (Devlin et al., 2019), we select top $k=20$ enriched terms for each class (c.f.
+Section 3.2).
+We generate $q=5$ documents with path-based generation for each class.
+The document encoder in the final classifier is initialized with BERT-base-uncased for a fair comparison with the baselines.
+We train the classifier using AdamW optimizer with a learning rate 5e-5, and the batch size is 64.
+The experiments are run on one NVIDIA RTX A6000 GPU.
+Table 3. Performance of TELEClass and its ablations on Amazon-531 and DBPedia-298 datasets. The best score is boldfaced.
 Methods Amazon-531 DBPedia-298
@@
 TELEClass 0.6483 0.8505 0.6421 0.6865 0.8633 0.9351 0.8633 0.8864
-Table 3: Performance comparison of TELEClass and zero-shot LLM prompting.
+4.2. Experimental Results
+Table 2 shows the evaluation results of all the compared methods.
+We make the following observations.
+(1) Overall, TELEClass achieves significantly better performance than other strong zero-shot and weakly-supervised baselines, which demonstrates the effectiveness of TELEClass on the hierarchical text classification task without any human supervision.
+(2) By comparing with other weakly-supervised methods, we find that TELEClass significantly outperforms TaxoClass-NoST, the strongest baseline that, like TELEClass, does not use self-training.
+Given that TELEClass uses an even simpler classifier model than TaxoClass-NoST, its superior performance shows the substantially better pseudo training data obtained by combining unlabeled corpus and LLMs.
+(3) Although LLMs (e.g., ChatGPT) show power in many tasks, naïvely prompting it in the hierarchical text classification task yields significantly inferior performance compared to strong weakly-supervised text classifiers.
+This proves the necessity of incorporating corpus-based knowledge to improve label taxonomy understanding for the hierarchical setting.
+We conduct a more detailed comparison with LLM prompting for hierarchical text classification in Section 4.4.
+We also study the temporal complexity of TELEClass.
+We observe that, on Amazon-531, both TELEClass and the strongest baseline TaxoClass take around 5 to 5.5 hours.
+The reason why TELEClass does not increase the overall temporal complexity is that TaxoClass needs to run the textual entailment model on each pair of document and candidate class.
+On the other hand, the taxonomy enrichment step of TELEClass makes it possible to simplify this process with embedding similarity calculation which saves a lot of time, while the saved time is budgeted for LLM prompting.
+Table 4.
+Performance comparison of TELEClass and zero-shot LLM prompting.
+We only report Example-F1 and P@k, because it is not straightforward to get ranking of classes predicted by LLMs for MRR calculation.
+We also report estimated costs in US dollars and running time in minutes for each method on the entire test set.
+“” indicates that we report the performance based on an estimation from a 1,000-document subset of test data.
 Methods Amazon-531 DBPedia-298
-F1 P@1 P@3 F1 P@1 P@3
-ChatGPT 0.516 0.681 0.475 0.482 0.533 0.455
-ChatGPT-L 0.662 0.857 0.644 0.665 0.830 0.649
-GPT-4 0.699 0.822 0.689 0.605 0.652 0.592
-TELEClass 0.648 0.851 0.642 0.863 0.935 0.863
-4.3 Ablation Studies
+Example-F1 P@1 P@3 Est. Cost Est. Time Example-F1 P@1 P@3 Est. Cost Est. Time
+GPT-3.5-turbo 0.5164 0.6807 0.4752 \$60 240 mins 0.4816 0.5328 0.4547 \$80 400 mins
+GPT-3.5-turbo (level) 0.6621 0.8574 0.6444 \$20 800 mins 0.6649 0.8301 0.6488 \$60 1,000 mins
+GPT-4 0.6994 0.8220 0.6890 \$800 400 mins 0.6054 0.6520 0.5920 \$2,500 1,000 mins
+TELEClass 0.6483 0.8505 0.6421 <\$1 3 mins 0.8633 0.9351 0.8633 <\$1 7 mins
+4.3. Ablation Studies
 We conduct ablation studies to better understand how each component of TELEClass contributes to final performance. Table 3 shows the results of the following ablations:
@@
 Even Gen-Only achieves comparable performance to the strong baseline TaxoClass-NoST, demonstrating the effectiveness of this augmentation step.
-4.4 Comparison with Zero-Shot LLM Prompting
+4.4. Comparison with Zero-Shot LLM Prompting
 In this section, we further compare TELEClass with zero-shot LLM prompting.
-Because it is not straightforward to get ranked predictions by LLMs, we only report Example-F1 and P@k as performance evaluations.
-We report the zero-shot performance of ChatGPT and GPT-4 by including all the classes in the prompt and asking the LLM to provide the three most appropriate classes for a given document.
-Experiments on GPT-4 are conducted on 1000 randomly sampled test documents given the prohibitively high cost.
-We also include a level-by-level version, ChatGPT-L, which selects one most appropriate class at each level and continues to its children.
+Because it is not straightforward to get ranked predictions by LLMs, we only report Example-F1 and P@k as performance evaluation.
+Additionally, we report the estimated cost and time for each method on the entire test set.
+The inference time is reported in minutes, and please be aware that this is just a rough estimation as the actual running time is also dependent on the server condition.
+We include the following settings to compare with TELEClass:
+  • GPT-3.5-turbo: We include all the classes in the prompt and ask GPT-3.5-turbo model to provide 3 most appropriate classes for a given document.
+  • GPT-3.5-turbo (level): We perform level-by-level prompting using GPT-3.5-turbo.
+Starting from the root node, we ask the model to return one most appropriate class for a given document, and we iteratively prompt the model with the children of the selected node at each level.
+This method can only generate a path in the taxonomy, but in the actual multi-label hierarchical classification setting, the true labels may not sit in the same path.
+  • GPT-4: We include all the classes in the prompt and ask GPT-4 to provide 3 most appropriate classes for a given document.
+Given the limited budget, we only test it on randomly sampled 1,000 documents and estimate the cost on the entire test set.
 Table 4 shows the experiment results.
-We find that TELEClass consistently outperforms zero shot performance of LLMs by large margins on DBPedia-298.
-On Amazon-531, while it continues to show substantial improvements over ChatGPT, it slightly underperforms ChatGPT-L and GPT-4.
-As for the cost, once trained, TELEClass does not require an additional cost on inference and also has a substantially shorter inference time.
-Prompting LLMs takes a longer time and can be prohibitively expensive, and the cost will scale up with an increasing size of test data.
-The estimated API cost for ChatGPT on the entire DBPedia test set is \$80 and for GPT-4 it is around \$2,500.
-Also, we find that the level-by-level version ChatGPT-L consistently outperforms the naïve version, demonstrating the necessity of taxonomy structure.
-Table 4: Intermediate results on two documents, including selected core classes by different methods, true labels, and the corresponding taxonomy enrichment results.
+We find that TELEClass consistently outperforms all compared methods on DBPedia, while on Amazon, TELEClass underperforms GPT-3.5-turbo (level) and GPT-4 but still being comparable.
+As for the cost, once trained, TELEClass does not require additional cost on inference and also has substantially shorter inference time.
+Prompting LLMs takes longer time and can be prohibitively expensive (e.g., using GPT-4), and the cost will scale up with increasing size of test data.
+Also, we find that GPT-3.5-turbo (level) consistently outperforms the naïve version, demonstrating the necessity of taxonomy structure.
+It saves the cost because of the much shorter prompts, but takes longer time due to more queries made per document.
+Table 5.
+Intermediate results on two documents, including selected core classes by different methods, true labels, and the corresponding taxonomy enrichment results.
 The optimal core class in the true labels is marked with ©.
 Dataset Document Core Classes by… True Labels Corr. Enrichment
-DBPedia The Lindenhurst Memorial Library (LML) is located in Lindenhurst, New York, and is one of the fifty six libraries that are part of the Suffolk Cooperative Library System … TaxoClass: village TELEClass initial: building TELEClass refined: library library©, agent, educational institution Class: library Enrichment: national library, central library, collection, volumes…
-Amazon Since mom (89 yrs young) isn’t steady on her feet, we have place these grab bars around the room.
+DBPedia The Lindenhurst Memorial Library (LML) is located in Lindenhurst, New York, and is one of the fifty six libraries that are part of the Suffolk Cooperative Library System … TaxoClass: village TELEClass initial: building TELEClass refined: library library©, agent, educational institution Class: library Top Enrichment: national library, central library, collection, volumes…
+Amazon Since mom (89 yrs young) isn’t steady on her feet, we have placed these grab bars around the room.
 It gives her the stability and security she needs.
-TaxoClass: personal care, health personal care, safety TELEClass initial: daily living aids, medical supplies equipment, safety, TELEClass refined: bathroom aids safety health personal care, medical supplies equipment, bathroom aids safety© Class: bathroom aids safety Enrichment: seat, toilet, shower, safety, handles…
-4.5 Case studies
-To better understand the TELEClass framework, we show some intermediate results of two documents in Table 4, including the core classes selected by (1) the original TaxoClass Shen et al.
-(2021) method, (2) TELEClass’s initial core classes selected by LLM, and (3) refined core classes of TELEClass.
+TaxoClass: personal care, health personal care, safety TELEClass initial: daily living aids, medical supplies equipment, safety, TELEClass refined: bathroom aids safety health personal care, medical supplies equipment, bathroom aids safety© Class: bathroom aids safety Top Enrichment: seat, toilet, shower, safety, handles…
+4.5. Case studies
+To better understand the TELEClass framework, we show some intermediate results of two documents in Table 5, including the core classes selected by (1) the original TaxoClass (Shen et al., 2021) method, (2) TELEClass’s initial core classes selected by LLM, and (3) refined core classes of TELEClass.
 Besides, we also include the true labels of the documents and the taxonomy enrichment results of the corresponding core class in the table.
@@
 In the other example, TELEClass also pinpoints the most accurate core class “bathroom aids safety” while other methods can only find more general or partially relevant classes.
-5 Related Work
-5.1 Weakly-Supervised Text Classification
-Weakly-supervised text classification trains a classifier with limited guidance, aiming to reduce human efforts while maintaining high proficiency.
-Various sources of weak supervision have been explored, including distant supervision Gabrilovich and Markovitch (2007); Chang et al.
-(2008); Song and Roth (2014) like knowledge bases, keywords Agichtein and Gravano (2000); Tao et al.
-(2018); Meng et al.
-(2018, 2020); Wang et al.
-(2021a); Zhang et al.
-(2021a), and heuristic rules Ratner et al.
-(2016); Badene et al.
-(2019); Shu et al.
-(2020).
+Besides, although TELEClass outperforms the zero-shot LLM prompting in most cases, there are cases showing the contrary and here is one example we found from Amazon-531.
+The product review is about “glycolic treatment pads” for which GPT correctly predicts its labels as “beauty” and “skin care”, while TELEClass predicts it as “health care”.
+We suspect that the word “treatment” in the review leads to the error because of the bias of term-based pseudo-labeling.
+It is a known issue of keyword-based methods and some solutions are proposed for the weakly-supervised flat text classification (Dong et al., 2023; Zhang et al., 2023a).
+We hope our study can motivate more research to solve this issue in the hierarchical setting.
+5. Related Work
+5.1. Weakly-Supervised Text Classification
+Weakly supervised text classification trains a classifier with limited guidance, aiming to reduce human efforts while maintaining high proficiency.
+Various sources of weak supervision have been explored, including distant supervision (Gabrilovich and Markovitch, 2007; Chang et al., 2008; Song and Roth, 2014) like knowledge bases, keywords (Agichtein and Gravano, 2000; Tao et al., 2018; Meng et al., 2018, 2020; Wang et al., 2021a; Zhang et al., 2021b), and heuristic rules (Ratner et al., 2016; Badene et al., 2019; Shu et al., 2020).
 Later, extremely weakly-supervised methods are proposed to solely rely on class names to generate pseudo labels and train classifiers.
-LOTClass Meng et al.
-(2020) utilizes MLM-based PLM as a knowledge base for extracting class-indicative keywords.
-X-Class Wang et al.
-(2021a) extracts keywords for creating static class representations through clustering.
-PIEClass Zhang et al.
-(2023c) employs PLMs’ zero-shot prompting to obtain pseudo labels with noise-robust iterative ensemble training.
-MEGClass Kargupta et al.
-(2023) acquires contextualized sentence representations to capture topical information at the document level.
-WOTClass Wang et al.
-(2023) ranks class-indicative keywords from generated classes and extracts classes with overlapping keywords.
-5.2 Hierarchical Text Classification
+LOTClass (Meng et al., 2020) utilizes MLM-based PLM as a knowledge base for extracting class-indicative keywords.
+X-Class (Wang et al., 2021a) extracts keywords for creating static class representations through clustering.
+WDDC (Zeng et al., 2022) prompts a masked language model and train a classifier over the word distributions.
+NPPrompt (Zhao et al., 2023) uses the PLM embeddings to construct verbalizers and their weighted distribution over each class.
+PESCO (Wang et al., 2023a) performs zero-shot classification with semantic matching between class description and then fine-tune a classifier with iterative contrastive learning.
+PIEClass (Zhang et al., 2023a) employs PLMs’ zero-shot prompting to obtain pseudo labels with noise-robust iterative ensemble training.
+MEGClass (Kargupta et al., 2023) acquires contextualized sentence representations to capture topical information at the document level.
+WOTClass (Wang et al., 2023b) ranks class-indicative keywords from generated classes and extracts classes with overlapping keywords.
+RulePrompt (Li et al., 2024) mutually enhances automatically mined logical rules and pseudo labels for fine-tuning a classifier.
+5.2. Hierarchical Text Classification
 A hierarchy provides a systematic top-to-down structure with inherent semantic relations that can assist in text classification.
 Typical hierarchical text classification can be categorized into two groups: local approaches and global approaches.
-Local approaches train multiple classifiers for each node or local structures Banerjee et al.
-(2019); Liu et al.
-(2005); Wehrmann et al.
-(2018).
-Global approaches, learn hierarchy structure into a single classifier through recursive regularization Gopal and Yang (2013), a graph neural network (GNN)-based encoder Peng et al.
-(2018); Huang et al.
-(2019); Zhou et al.
-(2020); Shen et al.
-(2021), or a joint document label embedding space Chen et al.
-(2021).
+Local approaches train multiple classifiers for each node or local structures (Banerjee et al., 2019; Liu et al., 2005; Wehrmann et al., 2018).
+Global approaches, learn hierarchy structure into a single classifier through recursive regularization (Gopal and Yang, 2013), a graph neural network (GNN)-based encoder (Peng et al., 2018; Huang et al., 2019; Zhou et al., 2020; Shen et al., 2021), or a joint document label embedding space (Chen et al., 2021).
+Recent studies also show that LLMs cannot comprehend the complex hierarchical structure (Bhambhoria et al., 2023; Fuhan et al., 2024; Yu et al., 2023a).
 Weak supervision is also studied for hierarchical text classification.
-WeSHClass Meng et al.
-(2019) uses a few keywords or example documents per class and pretrains classifiers with pseudo documents followed by self-training.
-TaxoClass Shen et al.
-(2021) follows the same setting as ours which uses the sole class name of each class as the only supervision.
+WeSHClass (Meng et al., 2019) uses a few keywords or example documents per class and pretrains classifiers with pseudo documents followed by self-training.
+TaxoClass (Shen et al., 2021) follows the same setting as ours which uses the sole class name of each class as the only supervision.
 It identifies core classes for each document using a textual entailment model, which is then used to train a multi-label classifier.
-Additionally, MATCH Zhang et al.
-(2021c) and HiMeCat Zhang et al.
-(2021b) study how to integrate associated metadata into the label hierarchy for document categorization with weak supervision.
-5.3 LLMs as Generators and Annotators
-Large language models (LLMs) have demonstrated impressive performance in many downstream tasks and are explored to help low-resource settings by synthesizing data as generators or annotators Ding et al.
-(2023).
-For data generation, few-shot examples Wang et al.
-(2021b) or class-conditioned prompts Meng et al.
-(2022) are explored for LLM generation and the generated data can be used as pseudo training data to further fine-tune a small model as the final classifier Ye et al.
-(2022).
+Additionally, MATCH (Zhang et al., 2021c) and HiMeCat (Zhang et al., 2021a) study how to integrate associated metadata into the label hierarchy for document categorization with weak supervision.
+5.3. LLMs as Generators and Annotators
+Large language models (LLMs) have demonstrated impressive performance in many downstream tasks and are explored to help low-resource settings by synthesizing data as generators or annotators (Ding et al., 2023).
+For data generation, few-shot examples (Wang et al., 2021b) or class-conditioned prompts (Meng et al., 2022) are explored for LLM generation and the generated data can be used as pseudo training data to further fine-tune a small model as the final classifier (Ye et al., 2022).
 Recently, Yu et al.
-(2023) proposed an attribute-aware topical text classification method that incorporates ChatGPT to generate topic-dependent attributes and topic-independent attributes to reduce topic ambiguity and increase topic diversity for generation.
-For data annotation, previous works utilize LLMs for unsupervised annotation Feng et al.
-(2021), Chain-of-Thought annotation with explanation generation He et al.
-(2024), and active annotation Zhang et al.
-(2023a).
-6 Conclusion and Future Work
+(2023b) proposed an attribute-aware topical text classification method that incorporates ChatGPT to generate topic-dependent attributes and topic-independent attributes to reduce topic ambiguity and increase topic diversity for generation.
+For data annotation, previous works utilize LLMs for unsupervised annotation (Feng et al., 2021), Chain-of-Thought annotation with explanation generation (He et al., 2024a), and active annotation (Zhang et al., 2023b).
+6. Conclusion and Future Work
 In this paper, we propose a new method, TELEClass, for the minimally-supervised hierarchical text classification task with two major contributions.
@@
 Second, we explore the utilization of LLMs in the hierarchical text classification in two directions: data annotation and data creation.
-On two public datasets, TELEClass can outperform existing baseline methods substantially, and we further demonstrate its effectiveness through ablation studies.
+On two public datasets, TELEClass can outperform existing baselines substantially, and we further demonstrate its effectiveness through ablation studies.
 We also conduct a comparative analysis of performance and cost for zero-shot LLM prompting for the hierarchical text classification task.
-For future works, first, we plan to generalize TELEClass’s idea into other low-resource text mining tasks with hierarchical label spaces, such as fine-grained entity typing.
-We also plan to explore how to extend TELEClass into more complicated label spaces, such as an extremely large space with millions of labels.
-Limitations
-In this paper, we propose TELEClass, a novel hierarchical text classification method that uses the label taxonomy as the only supervision.
-Despite its effectiveness shown in the experiments, there is still room to improve.
-First, TELEClass utilizes the general knowledge of pretrained LLM and shows strong performance on two general domain datasets.
-However, it remains a challenge to train effective text classifiers in private domains, for which LLMs lack relevant knowledge.
+For future works, first, we plan to generalize TELEClass’s idea of combining LLMs with data-specific knowledge into other low-resource text mining tasks with hierarchical label spaces, such as fine-grained entity typing.
 Second, in this paper, we mainly focus on acquiring high-quality pseudo labeled data while only utilizing the simplest classifier model and objective.
 It is worth studying how the proposed method can be further improved with more advanced network structure and noise-robust training objectives.
-Lastly, TELEClass assumes the availability of a large unlabeled corpus for pseudo labeling, whose performance is not fully studied if limited data is available.
-Ethics Statement
-In this work, we develop a new hierarchical text classification method and experiment on two publicly available datasets for research purposes.
-Overall, we do not foresee any significant ethical issues from our work.
-All the datasets are used as intended.
-We will open-source our code upon publication to assist future studies.
-Although showing strong language understanding capability, LLMs can sometimes generate misinformation Chen and Shu (2023) or biased contents Fang et al.
-(2024).
-In our work, instead of studying the generation ability of LLMs, we only use LLMs to generate pseudo documents for text classifier training.
-In practice, our method can be combined with any misinformation detection and debiasing methods Mattern et al.
-(2022); Chen and Shu (2021).
-Acknowledgments
+Lastly, we also plan to explore how to extend TELEClass into harder settings like when existing LLMs do not have the knowledge for the initial annotation (e.g., a private domain), a lower-resource scenario where the availability of the unlabeled corpus is limited, or a more complicated label space like an extremely large hierarchical label space with millions of classes.
+Acknowledgements.
 Research was supported in part by US DARPA INCAS Program No.
-HR001121C0165 and KAIROS Program No.
-FA8750-19-2-1004, National Science Foundation IIS-19-56151, and the Molecule Maker Lab Institute: An AI Research Institutes program supported by NSF under Award No.
+HR0011-21-C0165 and BRIES Program No.
+HR0011-24-3-0325, National Science Foundation IIS-19-56151, the Molecule Maker Lab Institute: An AI Research Institutes program supported by NSF under Award No.
 2019897, and the Institute for Geospatial Understanding through an Integrative Discovery Environment (I-GUIDE) by NSF under Award No.
@@
 Appendix A Pseudo code of TELEClass
-Algorithm 1 summarizes the entire TELEClass framework.
 Input: A corpus $\mathcal{D}$, a label taxonomy $\mathcal{T}$, a pretrained text encoder $\mathcal{S}$, an LLM $\mathcal{G}$.
@@
 13
-14      $T_{c}\leftarrow$ aggregate corpus-based and LLM-generated terms;
+14      $T_{c}\leftarrow$ aggregate corpus-based and LLM-generated terms Eq. 4;
 15
@@
 27 $\mathcal{D}^{\text{gen}}\leftarrow$ generate $q$ documents for each path using $\mathcal{G}$;
-28 $F\leftarrow$ train classifier with $\mathcal{D}^{\text{core}}$ and $\mathcal{D}^{\text{gen}}$;
+28 $F\leftarrow$ train classifier with $\mathcal{D}^{\text{core}}$ and $\mathcal{D}^{\text{gen}}$ Eq. 6;
 29 Return $F$;
 Algorithm 1 TELEClass
-Appendix B Experiment Setup
-B.1 Compared Methods
-We compare the following methods on the weakly-supervised hierarchical text classification task.
-  • Hier-0Shot-TC Yin et al. (2019) is a zero-shot approach, which utilizes a pretrained textual entailment model to iterative find the most similar class at each level for a document.
-  • Hier-doc2vec Le and Mikolov (2014) is a weakly-supervised approach, which first trains document and class representations in a same embedding space, and then iteratively selects the most similar class at each level.
-  • WeSHClass Meng et al.
-(2019) is a weakly-supervised approach using a set of keywords for each class.
-It first generates pseudo documents to pretrain text classifiers and then performs self-training.
-  • TaxoClass Shen et al.
-(2021) is a weakly-supervised approach that only uses the class name of each class.
-It first uses a textual entailment model with a top-down search and corpus-level comparison to select core classes, which are then used as pseudo training data.
-We include both its full model, TaxoClass, and its variation TaxoClass-NoST that does not apply self-training on the trained classifier, which is the same as TELEClass.
-B.2 Evaluation Metrics
-Let $\mathbb{C}^{\text{true}}_{i}$ and $\mathbb{C}^{\text{pred}}_{i}$ denote the set of true labels and the set of predicted labels for document $d_{i}\in\mathcal{D}$, respectively.
-If a method can generate rankings of classes within $\mathbb{C}^{\text{pred}}_{i}$, we further denote its top-k predicted labels as $\mathbb{C}^{\text{pred}}_{i,k}\subset\mathbb{C}^{\text{pred}}_{i}$.
-The evaluation metrics are defined as follows:
-  • Example-F1 Sørensen (1948), which is also called micro-Dice coefficient, evaluates the multi-label classification results without ranking,
-[…]
-  • Precision at k, or P@k, is a ranking-based metric that evaluates the precision of top-$k$ predicted classes,
-[…]
-  • Mean Reciprocal Rank, or MRR, is another ranking-based metric, which evaluates the multi-label predictions based on the inverse of true labels’ ranks within predicted classes,
-    $\text{MRR}=\frac{1}{|\mathcal{D}|}\sum_{d_{i}\in\mathcal{D}}\frac{1}{|\mathbb{C}_{i}^{\text{true}}|}\sum_{c_{j}\in\mathbb{C}_{i}^{\text{true}}}\frac{1}{\min\{k|c_{j}\in\mathbb{C}^{\text{pred}}_{i,k}\}}.$ (5)
-B.3 Implementation Details
-We use Sentence Transformer Reimers and Gurevych (2019) all-mpnet-base-v2 as the text encoder for the similarity measure in Section 3.1 and Section 3.3.
-We query GPT-3.5-turbo-0125 for LLM-based taxonomy enrichment, core class annotation, and path-based generation.
-For corpus-based taxonomy enrichment, we get the term and class name embeddings using a pre-trained BERT-base-uncased Devlin et al.
-(2019), we select top $k=20$ enriched terms for each class (c.f.
-Section 3.2).
-We generate $q=5$ documents with path-based generation for each class.
-The document encoder in the final classifier is initialized with BERT-base-uncased for a fair comparison with the baselines.
-We train the classifier using AdamW optimizer with a learning rate 5e-5, and the batch size is 64.
-The experiments are run on one NVIDIA RTX A6000 GPU.
-Appendix C Prompts for LLM
+Algorithm 1 summarizes the TELEClass method.
+Appendix B Prompts for LLM
   • LLM-based enrichment for Amazon-531
     Instruction: [Target Class] is a product class in Amazon and is the subclass of [Parent Class].
-Please generate 10 additional key terms about the [Target Class] that is relevant to [Target Class] but irrelevant to [Sibling Classes].
+Please generate 10 additional key terms about the [Target Class] that are relevant to [Target Class] but irrelevant to [Sibling Classes].
 Please split the additional key terms using commas.
   • LLM-based enrichment for DBPedia-298
-    Instruction: [Target Class] is a article category of Wikipedia articles and is the subclass of [Parent Class].
-Please generate 10 additional key terms about the [Target Class] that is relevant to [Target Class] but irrelevant to [Sibling Classes].
+    Instruction: [Target Class] is an article category of Wikipedia articles and is the subclass of [Parent Class].
+Please generate 10 additional key terms about the [Target Class] that are relevant to [Target Class] but irrelevant to [Sibling Classes].
 Please split the additional key terms using commas.
@@
     Instruction: You will be provided with an Amazon product review, and please select its product types from the following categories: [Candidate Classes].
+Just give the category names as shown in the provided list.
+Query: [Document]
+  • Core class annotation for DBPedia-298
+    Instruction: You will be provided with a Wikipedia article describing an entity at the beginning, and please select its types from the following categories: [Candidate Classes].
 Just give the category names as shown in the provided list.
@@
 You should focus on [The Leaf Node on the Path].
-  • Core class annotation for DBPedia-298
-    Instruction: You will be provided with a Wikipedia article describing an entity at the beginning, and please select its types from the following categories: [Candidate Classes].
-Just give the category names as shown in the provided list.
-Query: [Document]
   • Path-based generation for DBPedia-298
@@
 You should focus on [The Leaf Node on the Path].
+Appendix C Evaluation Metrics
+Let $\mathbb{C}^{\text{true}}_{i}$ and $\mathbb{C}^{\text{pred}}_{i}$ denote the set of true labels and the set of predicted labels for document $d_{i}\in\mathcal{D}$, respectively.
+If a method can generate rankings of classes within $\mathbb{C}^{\text{pred}}_{i}$, we further denote its top-k predicted labels as $\mathbb{C}^{\text{pred}}_{i,k}\subset\mathbb{C}^{\text{pred}}_{i}$.
+The evaluation metrics are defined as follows:
+  • Example-F1 (Sørensen, 1948), which is also called micro-Dice coefficient, evaluates the multi-label classification results without ranking,
+[…]
+  • Precision at k, or P@k, is a ranking-based metric that evaluates the precision of top-$k$ predicted classes,
+[…]
+  • Mean Reciprocal Rank, or MRR, is another ranking-based metric, which evaluates the multi-label predictions based on the inverse of true labels’ ranks within predicted classes,
+    (9) $\text{MRR}=\frac{1}{|\mathcal{D}|}\sum_{d_{i}\in\mathcal{D}}\frac{1}{|\mathbb{C}_{i}^{\text{true}}|}\sum_{c_{j}\in\mathbb{C}_{i}^{\text{true}}}\frac{1}{\min\{k|c_{j}\in\mathbb{C}^{\text{pred}}_{i,k}\}}.$